### v3

# Explicit estimates for the logarithmic derivative and the reciprocal of the Riemann zeta-function

Abstract
In this article, we improve on two explicit bounds of order $\log t$ for $\sigma$ close to $1$: $|\zeta^{\prime}(\sigma+it)/\zeta(\sigma+it)|$ and $|1/\zeta(\sigma+it)|$. Our method provides marked improvements in the constants, especially with regards to the latter. These savings come from replacing a lemma of Borel and Carathédory, with a Jensen-type one of Heath-Brown’s. Using an argument involving the trigonometric polynomial, we additionally provide a slight asymptotic improvement within the classical zero-free region: $1/\zeta(\sigma+it)\ll(\log t)^{11/12}$.
1 Introduction
In the study of the Riemann zeta function $\zeta(s)$ and related functions, one often requires upper bounds for the logarithmic derivative of $\zeta(s)$ and the reciprocal of $\zeta(s)$. For example, these come up in the Perron formula in the process of estimating the Chebyshev function $\psi(x)=\sum_{n\leq x}\Lambda(n)$, where $\Lambda(n)$ is the von Mangoldt function, or $M(x)=\sum_{n\leq x}\mu(n)$, where $\mu(n)$ is the Möbius function. The reader is referred to [27], [4], [9], [26, (3.14)] for more information on the topic.
Our main results will be presented in detail in Section 2, but here we highlight two of the more significant ones. In particular, Corollaries 4 and 5 of this article show respectively that for
$\sigma\geq 1-\frac{1}{5.5587\log t}\qquad\text{and}\qquad t\geq 1000,$
we have
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq 3.632\log t,\qquad\text{and}\qquad\left|\frac{1}{\zeta(\sigma+it)}\right|\leq 9.842(\log t)^{11/12}.$
Now we give a brief history of results on $\zeta^{\prime}(s)/\zeta(s)$ and $1/\zeta(s)$. Trudgian in [27] gives an explicit version of results in [26, §3] originally due to Landau, obtaining bounds of order $O(\log t)$. This was then extended in [14] and [7] to bounds of order $O(\log t/\log\log t)$. However, such estimates tend to have large constants, often rendering them unsuitable for practical use. It is a non-trivial task to reduce these constants. This is due to the fact that the factor of $\zeta(s)$ in the denominator means that such estimates can only be valid in a zero-free region, and the constants rapidly increase in size as we approach the boundary of the region, i.e., $\zeta(s)\to 0$.
The method in [26, §3] first estimates $|\zeta^{\prime}(s)/\zeta(s)|$, and uses the result to derive a bound for $|1/\zeta(s)|$. It relies on two applications of the Borel–Carathédory theorem. The first application serves to obtain a bound on ${\textup{Re}}(\zeta^{\prime}(s)/\zeta(s))$, which is then used in the second to arrive at the desired bound on $|\zeta^{\prime}(s)/\zeta(s)|$. As each application slightly worsens the final estimate, we were able to improve Trudgian’s results by cutting out the first application, and instead estimating ${\textup{Re}}(\zeta^{\prime}(s)/\zeta(s))$ using a lemma of Heath-Brown [13, Lemma 3.2].
While the improvements to bounds on $|\zeta^{\prime}(s)/\zeta(s)|$ given in this paper might seem minor compared to [27], we take the opportunity to point out an error in that paper. In the quantity $A$ of [27, Lemma 4], there should be a factor of $1/c$ instead of $1/2c$. This leads to the overall constants being better than they should be. For instance, in Table $2$ of said paper, which shows bounds of the form $|\zeta^{\prime}(s)/\zeta(s)|\leq R_{1}\log t$ for $\sigma\geq 1-1/W\log t$, one should have $(W,R_{1})=(6,473)$ and $(12,54.3)$, instead of $(6,382.58)$ and $(12,40.14)$, respectively. In this article, we have recovered the erroneous bounds and even slightly improved upon them.
Furthermore, we directly derive a bound on $|1/\zeta(s)|$ from ${\textup{Re}}(\zeta^{\prime}(s)/\zeta(s))$ rather than $|\zeta^{\prime}(s)/\zeta(s)|$, eliminating the need for the Borel–Carathédory theorem in this case, leading to a substantial improvement over [27]. In addition, when $\sigma=1$, the second assertion of Corollary 4 is significantly better than [3, Proposition A.2].
Finally, a novel argument utilising a combination of the classical non-negative trigonometric polynomial and a uniform bound of the form $\zeta(s)\ll(\log t)^{2/3}$ for $\sigma\geq 1$, allows us get an asymptotic improvement in the case of $1/\zeta(s)$. In particular, working within zero-free regions of the forms
$\sigma\geq 1-\frac{1}{Z_{1}\log t},\qquad\sigma\geq 1-\frac{\log\log t}{Z_{2}\log t},\qquad\sigma\geq 1-\frac{1}{Z_{3}(\log t)^{2/3}(\log\log t)^{1/3}},$
for some positive constants $Z_{1},\,Z_{2},\,Z_{3}$, one usually proves upper bounds for $1/\zeta(s)$ of the order
$\log t,\qquad\frac{\log t}{\log\log t},\qquad(\log t)^{2/3}(\log\log t)^{1/3},$ (1)
respectively. Without making any changes to the zero-free region used, our new method would allow one to refine the results in (1), to
$(\log t)^{11/12},\qquad\frac{(\log t)^{11/12}}{(\log\log t)^{3/4}},\qquad(\log t)^{2/3}(\log\log t)^{1/4},$ (2)
respectively. As far as we are aware, results of these types were not previously known. In this paper, we prove an explicit version of the first bound in (2). An explicit version of the third bound in (2) is given in [18].
2 Main results
Here we address matters of notation before listing some of the main results we prove.
2.1 Properties of the Riemann zeta function
[…]
The classical zero-free region for $\zeta(s)$ states that for $t\geq 2$,
$\zeta(s)\neq 0,\qquad\text{ for }\qquad\sigma\geq 1-\frac{1}{W_{0}\log t}.$
Currently, the best known result is due to [20], where above $H_{0}$, the value $W_{0}=5.558691$ is permissible. We take $W_{0}$ to be this value.
We denote the Riemann hypothesis up to height $T$ by $RH(T)$, by which we mean: for all non-trivial zeros $\rho=\beta+i\gamma$ such that $\zeta(\rho)=0$ and $\gamma\leq T$, we have $\beta=1/2$. Similarly we denote the Riemann hypothesis by $RH$, by which we mean all non-trivial zeros $\rho$ have $\beta=1/2$.
2.2 Bounds on $|\zeta^{\prime}(s)/\zeta(s)|$
Here we present our main results, which are corollaries of the theorems in Sections 4 and 5.
Corollary 1.
Let $W>W_{0}$. In a region
$\sigma\geq 1-\frac{1}{W\log t}$
where $\zeta(\sigma+it)\neq 0$, we have for $t\geq 5$,
$\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq Q\log t,$
where $(W,Q)=(10,49.359)$ is admissible, and other values of $Q$ with corresponding $W$ are presented in Table 3.
Corollary 2.
For $\sigma\geq 1$ and $t\geq 5$, we have
$\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq 18.062\log t.$
Corollary 3.
Assume $RH(H_{0})$. Then in the region
$\sigma\geq\sigma_{0}>1/2\qquad\text{ and }\qquad 5\leq t\leq H,$
we have
$\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq Q_{RH(H_{0})}\log t,$
where $(\sigma_{0},Q_{RH(H_{0})})=(0.8,33.249)$ is admissible, and other values for $Q_{RH(H_{0})}$ with corresponding values of $\sigma_{0}$ are given in Table 2. Furthermore, by assuming $RH$ instead of $RH(H_{0})$, the upper bound restriction on $t$ can be removed.
2.3 Bounds on $|1/\zeta(s)|$
Corollary 4.
Let $W>W_{0}$. In a region
$\sigma\geq 1-\frac{1}{W\log t}$
where $\zeta(\sigma+it)\neq 0$, we have for $t\geq t_{0}$,
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq Y(W,t_{0})\log t,$
where $Y(5.5587,10^{3})=3.632$ is admissible, and other values of $Y(W_{0},t_{0})$ with corresponding $t_{0}$ are presented in Table 5, and values of $Y(W,5)$ with corresponding $W$ are presented in Table 6.
In addition, we have for $\sigma=1$ and $t\geq 2$,
$\left|\frac{1}{\zeta(1+it)}\right|\leq 2.751\log t.$
Corollary 5.
Let $W>W_{0}$. In a region
$\sigma\geq 1-\frac{1}{W\log t}$
where $\zeta(\sigma+it)\neq 0$, we have for $t\geq t_{0}$,
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq Y^{\prime}(W,t_{0})(\log t)^{11/12},$
where $Y^{\prime}(5.5587,10^{3})=9.842$ is admissible, and other values of $Y^{\prime}(W,t_{0})$ with corresponding $W$ and $t_{0}$ are presented in Table 7.
In addition, we have for $\sigma=1$ and $t\geq 2$,
$\left|\frac{1}{\zeta(1+it)}\right|\leq 7.475(\log t)^{11/12}.$
3 Some preliminaries
In this section, we present the results used in obtaining our estimates. For convenience, we begin by stating a few function theoretic results, which lie at the heart of our method.
Theorem 1.
[14, Theorem 4][Borel–Carathéodory] Let $s_{0}$ be a complex number, and let $R$ be a positive number possibly dependent on $s_{0}$. Suppose that the function $f(s)$ is analytic in a region containing the disc $|s-s_{0}|\leq R$. Let $M$ denote the maximum of $|\textup{Re}\,f(s)|$ on the boundary $|s-s_{0}|=R$. Then, for any $r\in(0,R)$ and any $s$ such that $|s-s_{0}|\leq r$,
$|f(s)|\leq\frac{2r}{R-r}M+\frac{R+r}{R-r}|f(s_{0})|.$
Seeking to apply Theorem 1 with $f(s)=-\zeta(s)^{\prime}/\zeta(s)$, we obtain a bound on $M$ using the following lemma due to Heath-Brown [13, Lemma 3.2]. The lemma is related to the well-known one of Jensen’s, and also to one of Carleman’s (see [25, §3.71]).
Lemma 1.
[13, Lemma 3.2] Let $f(z)$ be holomorphic for $|z-a|\leq R$, and non-vanishing both at $z=a$ and on the circle $|z-a|=R$. Let $\rho_{j}=a+r_{j}\exp(i\theta_{j})$ be the zeros of $f(z)$ in the disc, where $\rho_{j}$ has multiplicity $n_{j}$. Then
${\textup{Re}}\frac{f^{\prime}}{f}(a)=-\sum_{\rho_{j}}n_{j}\left(\frac{1}{r_{j}}-\frac{r_{j}}{R^{2}}\right)\cos\theta_{j}+\frac{1}{\pi R}\int_{0}^{2\pi}(\cos\theta)\log|f(a+Re^{i\theta})|\text{d}\theta.$
Lemma 2.
Let $d>0$ be some real constant and suppose that $\zeta(s)\neq 0$ in the region $\sigma\geq 1-1/W_{0}\log t$. Let $R^{\prime}>0$ be a parameter, chosen such that
$\frac{d}{\log t}\leq\frac{1}{W_{0}\log(t+R^{\prime})}.$ (3)
Then for $t\geq 3$, we have
$-{\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(\sigma+it)\leq\frac{4}{\pi R^{\prime}}\max_{0\leq\theta\leq 2\pi}|\log|\zeta(\sigma+it+R^{\prime}e^{i\theta})||,$
which holds uniformly for
$\sigma\geq 1-\frac{d}{\log t}.$ (4)
Proof.
For $\sigma_{0}$ in the region stated in (4), we take $R=R^{\prime}$, $f(s)=\zeta(s)$, $a=s_{0}=\sigma_{0}+it_{0}$, and $\rho_{k}=s_{0}+r_{k}\exp(i\theta_{k})$ to be a zero of $\zeta(s)$ in Lemma 1 (where we will make the substitution $s_{0}\to s$ at the end of the proof). This gives us
${\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(s_{0})=-\sum_{\rho_{k}}n_{k}\left(\frac{1}{r_{k}}-\frac{r_{k}}{R^{\prime 2}}\right)\cos\theta_{k}+\frac{1}{\pi R^{\prime}}\int_{0}^{2\pi}(\cos\theta)\log|\zeta(s_{0}+R^{\prime}e^{i\theta})|\text{d}\theta.$
Note that if a zero of $\zeta(s)$ lies on the boundary $|s-s_{0}|=R^{\prime}$, then since the zeros are separable, we can always find an arbitrarily smaller $R^{\prime}$ where the boundary is free of zeros.
For convenience, we define
$J:=\int_{0}^{2\pi}(\cos\theta)\log|\zeta(s_{0}+R^{\prime}e^{i\theta})|\text{d}\theta,$
for which we seek a lower bound.
For $20\leq\theta\leq\pi/$ 2 and $3\pi/2\leq\theta\leq 2\pi$, i.e. when $0\leq\cos\theta\leq 1$, we have
$\log|\zeta(s_{0}+R^{\prime}e^{i\theta})|\geq-\max_{0\leq\theta\leq 2\pi}|\log|\zeta(s_{0}+R^{\prime}e^{i\theta})||,$
so in the ranges $20\leq\theta\leq\pi/$ 2 and $3\pi/2\leq\theta\leq 2\pi$ combined,
$J\geq-2\max_{0\leq\theta\leq 2\pi}|\log|\zeta(s_{0}+R^{\prime}e^{i\theta})||.$
For the remaining range $\pi/2\leq\theta\leq 3\pi/2$, noting that $-1\leq\cos\theta\leq 0$, we have
$\int_{\pi/2}^{3\pi/2}(\cos\theta)$ $\log|\zeta(s_{0}+R^{\prime}e^{i\theta})|\text{d}\theta\geq-2\max_{0\leq\theta\leq 2\pi}|\log|\zeta(s_{0}+R^{\prime}e^{i\theta})||,$
and hence for the whole range $\theta\in[0,2\pi]$,
$J\geq-4\max_{0\leq\theta\leq 2\pi}|\log|\zeta(s_{0}+R^{\prime}e^{i\theta})||.$
Now Lemma 1 yields
$-{\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(s_{0})=\sum_{|s_{0}-\rho|\leq R^{\prime}}\left(\frac{1}{r_{k}}-\frac{r_{k}}{R^{\prime 2}}\right)\cos\theta_{k}-\frac{1}{\pi R^{\prime}}J,$ (5)
and since $r_{k}\cos\theta_{k}=-{\textup{Re}}(s_{0}-\rho_{k})$ and
$\frac{\cos\theta_{k}}{r_{k}}=\frac{r_{k}\cos\theta_{k}}{r_{k}^{2}}=\frac{-{\textup{Re}}(s_{0}-\rho_{k})}{|s_{0}-\rho_{k}|^{2}}=-{\textup{Re}}\left(\frac{1}{s_{0}-\rho_{k}}\right),$
the sum over all zeros in $|s_{0}-\rho|\leq R^{\prime}$ from (5) is
$\sum_{|s_{0}-\rho|\leq R^{\prime}}\left(\frac{1}{r_{k}}-\frac{r_{k}}{R^{\prime 2}}\right)\cos\theta_{k}$ $=\sum_{|s_{0}-\rho|\leq R^{\prime}}-{\textup{Re}}\left(\frac{1}{s_{0}-\rho}-\frac{s_{0}-\rho}{R^{\prime 2}}\right)$
$=-\sum_{|s_{0}-\rho|\leq R^{\prime}}(\sigma_{0}-\beta)\left(\frac{1}{|s_{0}-\rho|^{2}}-\frac{1}{R^{\prime 2}}\right),$ (6)
where the second factor in (3) is non-negative.
Notice that if
$1-\frac{1}{W_{0}\log(t_{0}+R^{\prime})}\leq\sigma_{0},$
then by our non-vanishing condition on $\zeta(s)$, we always have that $\sigma_{0}-\beta\geq 0$ in the disc $|s_{0}-\rho|\leq R^{\prime}$. This condition is fulfilled by our choice of $R^{\prime}$ in (3). Thus the entire right-hand side of (3) is negative, and so we can discard any contribution from this.
Combining this with the bound on $J$ and (5) proves the lemma. ∎
The following corollary is useful for proving results in a right-half plane, under $RH(T)$ or $RH$.
Corollary 6.
Let $\epsilon>0$ be a constant and suppose that $\zeta(s)\neq 0$ in the region $\sigma>1/2$ and $t\leq t+\epsilon$. Then for a fixed $R^{\prime}<\epsilon$ and $t\geq 3$, we have
$-{\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(\sigma+it)\leq\frac{4}{\pi R^{\prime}}\max_{0\leq\theta\leq 2\pi}|\log|\zeta(\sigma+it+R^{\prime}e^{i\theta})||,$
which holds uniformly for
$\sigma\geq\frac{1}{2}+\epsilon.$ (7)
Proof.
The proof is analogous to that of Lemma 2 with the additional assumption that the disc $|s_{0}-\rho|\leq R^{\prime}$ is completely free of zeros. ∎
Next we require some uniform bounds for $\zeta(s)$. Lemma 4 slightly improves [28, Corollary 1], which is proven using the Phragmén–Lindelöf Principle (Lemma 3).
Lemma 3.
[29, Lemma 3] Let $a,b,Q$ be real numbers such that $b>a$ and $Q+a>1$. Let $f(s)$ be a holomorphic function on the strip $a\leq\textup{Re}(s)\leq b$ such that
$|f(s)|<C\exp(e^{k|t|})$
for some $C>0$ and $0<k<\tfrac{\pi}{b-a}$. Suppose further that there are $A,B,\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}\geq 0$ such that $\alpha_{1}\geq\beta_{1}$ and
$|f(s)|\leq\begin{cases}A|Q+s|^{\alpha_{1}}\log^{\alpha_{2}}|Q+s|\quad\text{for}\textup{Re}(s)=a;\\ B|Q+s|^{\beta_{1}}\log^{\beta_{2}}|Q+s|\quad\text{for }\textup{Re}(s)=b.\end{cases}$
Then for $a\leq\textup{Re}(s)\leq b$, one has
$|f(s)|\leq\Big{(}{A|Q+s|^{\alpha_{1}}\log^{\alpha_{2}}|Q+s|}\Big{)}^{\tfrac{b-\textup{Re}(s)}{b-a}}\Big{(}{B|Q+s|^{\beta_{1}}\log^{\beta_{2}}|Q+s|}\Big{)}^{\tfrac{\textup{Re}(s)-a}{b-a}}.$
Lemma 4.
Let there be a real number $h(s)=(s-1)\zeta(s)$, and let $\delta_{r}$ be a positive real number. Furthermore, let $Q_{0}\geq 0$ be a number for which
$|h(\tfrac{1}{2}+it)|$ $\leq k_{1}|Q_{0}+\tfrac{1}{2}+it|^{k_{2}+1}(\log|Q_{0}+\tfrac{1}{2}+it|)^{k_{3}}$ (8)
$|h(1+\delta_{r}+it)|$ $\leq\zeta(1+\delta_{r})|Q_{0}+1+\delta_{r}+it|,$
for all $t$. Then for $\sigma\in[\tfrac{1}{2},1+\delta_{r}]$ and $t\geq t_{0}\geq e$, we have
$|\zeta$ $(s)|\leq\left(\zeta(1+\delta_{r})(1+a_{1}(1+\delta_{r},Q_{0},t_{0}))\right)^{\frac{\sigma-0.5}{0.5+\delta_{r}}}$ (9)
×$\bm{\times}\left(k_{1}(1+a_{0}(1+\delta_{r},Q_{0},t_{0}))^{k_{3}}(1+a_{1}(1+\delta_{r},Q_{0},t_{0}))^{k_{2}+1}t^{k_{2}}\log^{k_{3}}t\right)^{\frac{1+\delta_{r}-\sigma}{0.5+\delta_{r}}},$
where
$a_{0}(\sigma,Q_{0},t)=\frac{\sigma+Q_{0}}{2t^{2}\log t}+\frac{\pi}{2\log t}+\frac{\pi(\sigma+Q_{0})^{2}}{4t\log^{2}t},\qquad a_{1}(\sigma,Q_{0},t)=\frac{\sigma+Q_{0}}{t},$
and the right-hand side of (9) is decreasing in $\sigma$, provided that $k_{i}\geq 0$, and the following hold:
$t^{k_{2}}\log^{k_{3}}t\geq\frac{k_{1}(1+a_{0})^{k_{3}}(1+a_{1})^{k_{2}}}{\zeta(1+\delta_{r})}.$ (10)
Proof.
The proof is analogous to that of [28, Lemma 2.1], except that we are more precise in estimating (9) by retaining the dependence on $\sigma$ in the exponents of $1+a_{1}$ and $1+a_{0}$.
Note that different to the proof of [28, Lemma 2.1], we use only two lines, $\sigma=1/2$ and $\sigma=1+\delta_{r}$, omitting the need for a bound at $\sigma=1$. This is because we intend to find the maximum bound over $\sigma\in[1/2,1+\delta_{r}]$, which we show occurs at $\sigma=1/2$, as our bound is decreasing (by the condition (10)) in $\sigma$. Additionally, we found that further splitting the interval complicates (10) and raises the threshold of $t$ for the lemma to be valid. ∎
Corollary 7.
For $\sigma\in[\tfrac{1}{2},1+\delta_{r}]$ and $t\geq t_{0}\geq e$, we have
$|\zeta(s)|\leq 0.618(1+a_{0}(1+\delta_{r},1.31,t_{0}))(1+a_{1}(1+\delta_{r},1.31,t_{0}))^{7/6}t^{1/6}\log t,$
provided that
$t^{1/6}\log t\geq\frac{0.618(1+a_{0})(1+a_{1})^{1/6}}{\zeta(1+\delta_{r})}.$ (11)
Proof.
Recall that Hiary, Patel, and Yang [15] have $|\zeta(1/2+it)|\leq 0.618t^{1/6}\log t$, for $t\geq 3$.
We now refer to Lemma 4, taking the maximum of (9) at $\sigma=1/2$. It has been shown in [30, p. 15] that $Q_{0}=1.31$ is permissible in (8) together with
$(k_{1},k_{2},k_{3},Q_{0})=\left(0.618,\tfrac{1}{6},1,1.31\right),$
hence we are done. ∎
Lemma 5.
For every $|t|\geq 3$, we have
$|\zeta(1+it)|\leq 58.096(\log|t|)^{2/3}.$
Proof.
We use same argument in the first paragraph of [29, p. 2], which modifies Ford’s proof in [10]. The only difference is that we now apply the argument to [2, §5], which itself is an improvement on Ford’s result. This proves the lemma. ∎
Lemma 6.
For $t\geq t_{0}\geq 3$ and $\sigma\geq 1$, we have
$|\zeta(\sigma+it)|\leq C_{3}(\log t)^{2/3},$
with
$C_{3}=58.096\sqrt{1+9t_{0}^{-2}}\left(1+\frac{\log\sqrt{1+9t_{0}^{-2}}}{\log t_{0}}\right).$ (12)
Proof.
The proof is similar to that of [14, Lemma 7]. First, to show that the result holds for $1\leq\sigma\leq 2$, we use the Phragmén–Lindelöf Principle, i.e., Lemma 3, on the holomorphic function
$f(s)=(s-1)\zeta(s).$
To this end, on the $1$-line we have
$|f(1+it)|=|t|\,|\zeta(1+it)|\leq 58.096\,|2+it|(\log|2+it|)^{2/3}.$
This inequality is verified numerically for $|t|<3$, and is a consequence of Lemma 5 for $|t|\geq 3$.
Also, on the line $s=2+it$,
$|f(2+it)|\leq|1+it|\,\zeta(2)<58.096|3+it|(\log|3+it|)^{2/3},$
for all real $t$. Now apply Lemma 3 with the following parameter values: $a=1$, $b=2$, $Q=1$, $\alpha_{1}=\beta_{1}=1$, $\alpha_{2}=\beta_{2}=2/3$, and $A=B=58.096$. We obtain
$|f(s)|\leq 58.096|1+s|(\log|1+s|)^{2/3},$ (13)
for $1\leq\sigma\leq 2$. Moreover, for the same range of $\sigma$, we have the following estimates:
$\begin{split}\frac{|1+s|}{|s-1|}&\leq\frac{\sqrt{9+t^{2}}}{t}\leq\sqrt{1+9t_{0}^{-2}},\\ \log|1+s|&\leq\log\sqrt{9+t^{2}}\leq\log t+\log\sqrt{1+9t_{0}^{-2}}.\end{split}$
Thus for $1\leq\sigma\leq 2$ and $t\geq t_{0}$, the inequality (13) implies that
$|\zeta(\sigma+it)|$ $\leq 58.096\sqrt{1+9t_{0}^{-2}}\left(1+\frac{\log\sqrt{1+9t_{0}^{-2}}}{\log t_{0}}\right)\log^{2/3}t<C_{3}\log^{2/3}t,$
as desired, where $C_{3}$ is defined in (12).
Finally, we address the case where $\sigma\geq 2$. Here since $\sigma>1$, we apply the trivial bound in (49). Then for $\sigma\geq 2$, it is easy to verify that
$|\zeta(\sigma+it)|\leq 2\leq C_{3}(\log t)^{2/3}$
for all $t\geq t_{0}$. ∎
Lemma 7.
[23, Lemma 4] In the region 1/$21/2<\sigma\leq 3/$ 2 and $0\leq t\leq 2\exp(e^{2})$, we have
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq\frac{4}{\sigma-(1/2)}.$
In the remainder of this section, we present some estimates concerning the Riemann zeta function, including some for $\sigma>1$, which are better than the trivial ones. We would like to point out that Lemma 8 and Corollary 9, which can both be viewed as an extension of Corollary 8 inside the critical strip, are not actually required for the purposes of this paper. However, we include them here as the proofs are similar to that of Corollary 8, and at the same time we provide a brief discussion on their usefulness. Lemma 8 is an improvement on [16, Proposition 5.3.2], which proves a similar result.
Lemma 8.
Let $W>0$ be a real constant. Then in the region
$1-\frac{1}{W\log t}\leq\sigma\leq\sigma_{1},$
we have for $t\geq t_{0}\geq 3$ and any $\eta\in[2/t_{0},1-1/t_{0}]$ that
$|\zeta(\sigma+it)|\leq e^{1/W}\log t+C_{0}(W,\sigma_{1},t_{0},\eta)$
where
$C_{0}(W,\sigma_{1},t_{0},\eta)=$ $e^{1/W}\Bigg{(}\log\left(\eta+\frac{1}{t_{0}}\right)+\gamma+\frac{1}{12\eta^{2}}$
$+\frac{1}{t_{0}}\left(\frac{1}{6\eta^{2}}+\frac{1}{\eta}+1\right)+\frac{16(\sigma_{1}^{2}+2\sigma_{1}-1)+3}{192\eta^{2}t_{0}^{2}}\Bigg{)}.$
Proof.
We will use the following representation for $\zeta(s)$ for $N\geq 2$,
$\zeta(s)=\sum_{n=1}^{N-1}\frac{1}{n^{s}}+\frac{1}{2N^{s}}+\frac{N^{1-s}}{s-1}+\frac{s}{12N^{s+1}}-\frac{s(s+1)}{2}\int_{N}^{\infty}\frac{\varphi^{*}(u)}{u^{s+2}}\text{d}u,$ (14)
where $\varphi^{*}(u)$ is the periodic function obtained by extending the polynomial $\varphi(u)=u^{2}-u+1/6$ on the interval $[0,1]$. Note that (14) holds for $\sigma>0$ by analytic continuation (see [26, Theorem 3.5]).
In the region of $\sigma$ considered, we have
$|N^{-s}|=N^{-\sigma}\leq\exp\left(-\Big{(}1-\frac{1}{W\log t}\Big{)}\log N\right)=N^{-1}\exp\left(\frac{\log N}{W\log t}\right).$ (15)
This bound can be used throughout (14). The first term in (14) can also be bounded using the estimate in [11, Lemma 2.1], so for $N\geq 2$,
$\left|\sum_{n=1}^{N-1}\frac{1}{n^{s}}+\frac{1}{2N^{s}}\right|$ $=\left|\sum_{n=1}^{N}\frac{1}{n^{s}}-\frac{1}{2N^{s}}\right|$
$\leq\exp\left(\frac{\log N}{W\log t}\right)\left(\sum_{n=1}^{N}\frac{1}{n}+\frac{1}{2N}\right)$
$\leq\exp\left(\frac{\log N}{W\log t}\right)\left(\log N+\gamma+\frac{1}{N}-\frac{1}{12N^{2}}+\frac{1}{64N^{4}}\right).$
The term $N^{1-s}/(s-1)$ will be bounded by $t^{-1}\exp\left(\frac{\log N}{W\log t}\right)$. The integral in (14) can be evaluated using the bound $|\varphi^{*}(u)|\leq 1/6$, so
$\int_{N}^{\infty}\frac{|\varphi^{*}(u)|}{u^{\sigma+2}}\text{d}u\leq\frac{1}{6(\sigma+1)}\frac{1}{N^{\sigma+1}}\leq\frac{1}{6}\exp\left(\frac{\log N}{W\log t}\right)\frac{1}{N^{2}}.$
This implies
$\left|\frac{s}{12N^{s+1}}-\frac{s^{2}+s}{2}\int_{N}^{\infty}\frac{\varphi^{*}(u)}{u^{s+2}}\text{d}u\right|$ $\leq\exp\left(\frac{\log N}{W\log t}\right)\left(\frac{2|s|}{12N^{2}}+\frac{|s^{2}|}{12N^{2}}\right)$
$\leq\exp\left(\frac{\log N}{W\log t}\right)\left(\frac{t+\sigma_{1}}{6N^{2}}+\frac{t^{2}+\sigma_{1}^{2}}{12N^{2}}\right).$
We thus have that $|\zeta(s)|$ is upper bounded by
$\exp\left(\frac{\log N}{W\log t}\right)\left(\log N+\gamma+\frac{1}{N}+\frac{1}{64N^{4}}+\frac{1}{t}+\frac{t^{2}+\sigma_{1}^{2}+2(t+\sigma_{1})-1}{12N^{2}}\right).$
It remains to choose $N$. We can minimise the expression by taking $N=\lfloor\eta t\rfloor+1$ for some constant $\eta\in[2/t_{0},1-1/t_{0}]$. This will make
$\exp\left(\frac{\log N}{W\log t}\right)\leq\exp\left(\frac{1}{W}\right)$
for all $t\geq t_{0}$. The bound then becomes
$|\zeta(s)|$ $\leq e^{1/W}\left(\log t+\log(\eta+t^{-1})+\gamma+\frac{1}{t}\left(\frac{1}{\eta}+1\right)+\frac{1}{64\eta^{2}t^{2}}\right.$
$\hskip 142.26378pt+\left.\frac{1}{12\eta^{2}}+\frac{1}{6\eta^{2}t}+\frac{\sigma_{1}^{2}+2\sigma_{1}-1}{12\eta^{2}t^{2}}\right)$
$\leq e^{1/W}\log t+C_{0}(W,\sigma_{1},t_{0},\eta)$
for $t\geq t_{0}$. ∎
We also provide a corollary, specialised for $\sigma\geq 1$, which sharpens the above estimates.
Corollary 8.
Let $s=\sigma+ikt$ with $1\leq\sigma\leq\sigma_{1}\leq 2$, $t\geq t_{0}\geq 3$, and a fixed $k\geq 1$. For $0<\eta/k<t_{0}$, we have
$|\zeta(\sigma+ikt)|<\log kt+C(\sigma_{1},t_{0},k,\eta),$
where
$C(\sigma_{1},t_{0},k,\eta)=$ $\log\left(\frac{1}{\eta}+\frac{1}{kt_{0}}\right)+\gamma+\frac{\eta^{4}}{64(kt_{0})^{4}}$ (16)
$+\frac{12+(c_{1}(\sigma_{1},t_{0},k)-(kt_{0})^{-1})\eta^{2}}{12kt_{0}}+\frac{c_{2}(\sigma_{1},t_{0},k)\eta^{2}}{24},$
with $\gamma$ denoting Euler’s constant, and
$c_{1}(\sigma_{1},t_{0},k)$ $=\left(\frac{\sigma_{1}^{2}}{(kt_{0})^{2}}+1\right)^{1/2},$
$c_{2}(\sigma_{1},t_{0},k)$ $=c_{1}(\sigma_{1},t_{0},k)\left(\frac{(\sigma_{1}+1)^{2}}{(kt_{0})^{2}}+1\right)^{1/2}.$
Proof.
The proof follows the argument of [3, Proposition A.1], originally due to Backlund [1], and extends their result to $t<500$. We omit the proof since it is almost identical to that of Lemma 8. ∎
To use Lemma 8, we want to choose a $t_{0}$ and optimise over $\eta$. For the smallest $t_{0}=3$, we can only use $\eta=2/3$. However, as $t_{0}\rightarrow\infty$, the optimal $\eta$ quickly approaches $1/\sqrt{6}\approx 0.4082$. We also have the following corollary, which is not required for our purposes, but is useful to have for other applications.
Corollary 9.
In the region
$1-\frac{1}{W_{0}\log t}\leq\sigma\leq 1,$ (17)
we have for $t\geq t_{0}\geq 3$,
$|\zeta(s)|\leq 1.1971\log t+B_{1},$ (18)
where Table 1 gives values of $B_{1}$ for select $t_{0}$.
Proof.
For a fixed $t_{0}$ and $\sigma_{1}=1$, in Lemma 8 optimise over $\eta$ to find the minimal $B_{1}$. ∎
      $t_{0}$       $B_{1}$       $\eta$
      $3$       $2.1173$       $2/3$
      $10^{3}$       $0.2254$       $0.41$
      $10^{12}$       $0.2171$       $0.41$
Table 1: Values of $B_{1}$ for Corollary 9 with the corresponding optimised $\eta$ from Lemma 8, where $\sigma_{1}=1$.
Corollary 9 can be used in conjunction with other estimates for $\zeta(s)$ in the critical strip, such as Ford’s Theorem 1 in [10] or Trudgian’s Lemma 2.1 [28]. Ford’s bound, while asymptotically better, is only sharper than bounds of the form $\zeta(s)\ll\log t$ for very large $t$. For $\sigma\in[1-1/(W_{0}\log t),1]$ and $3\leq t\leq 10^{10^{5}}$, Corollary 9 is an improvement to [10, Theorem 1]. This makes it useful for results over smaller $t$.
It is not uncommon to degrade asymptotically sharper results for $\zeta(s)$ to $|\zeta(s)|\leq C\log t$ for the sake of a smaller main-term constant $C$. Corollary 9 now directly offers a better constant. For instance, over $\sigma\in[1-1/(W_{0}\log t),1]$ and $t\geq 3$ our result implies
$|\zeta(s)|\leq 3.13\log t,$
whereas degrading a Richert-type bound like Ford’s [10, Theorem 1] would give
$|\zeta(s)|\leq 102\log t,$
and a bound obtained via the Phragmén–Lindelöf principle like (2.4) in [28, Lem. 2.1] using estimates on the $1/2$-line (see Hiary, Patel, and Yang [15]) and estimates on the $1$-line (see Backlund [1]) would yield
$|\zeta(s)|\leq 364\log t.$
We also remark that although Lemma 6 is asymptotically superior to Corollary 8, one would expect that the former beats the latter only for a sufficiently large $t$; approximately $t\geq\exp(58^{3})\geq 10^{84736}$, say. This can be improved by lowering the constant in Lemma 5, and correspondingly $C_{3}$.
Next we have some bounds on $\zeta(\sigma)$ for real $\sigma>1$. These are useful in their own right, but we show they can be used to obtain good bounds on $-\zeta^{\prime}(\sigma)/\zeta(\sigma)$ in a similar region.
Lemma 9.
[22, Lemma 5.4] Let $\gamma$ denote Euler’s constant. For $\sigma>1$ we have
$|\zeta(\sigma+it)|\leq\zeta(\sigma)\leq\frac{e^{\gamma(\sigma-1)}}{\sigma-1}.$
Lemma 10.
Let $\gamma_{n}$ be the Laurent–Stieltjes constants (with $\gamma:=\gamma_{0}$). For integers $k$ such that $1\leq k\leq 1500$, and real $\sigma$ satisfying $31<\sigma\leq\sigma_{k}<$ 3, we have
$\zeta(\sigma)\geq\frac{1}{\sigma-1}+\gamma+\phi_{0}(\sigma,k),$
where
$\phi_{0}(\sigma,k)=\sum_{n=1}^{k}\frac{(-1)^{n}(\sigma-1)^{n}}{n!}\gamma_{n}+\frac{1}{(\sigma-3)}\left(\frac{\sigma-1}{2}\right)^{k+1}\geq 0.$ (19)
In addition, we note some admissible pairs:
$(\sigma_{k},k)=(1.45,1),\,(2,3),\,(2.51,10),\,(2.97,500),\,(2.98,1000),\,(2.99,1500).$
Proof.
Recall from the Laurent series expansion of the zeta function
$\zeta(\sigma)=\frac{1}{\sigma-1}+\gamma+\sum_{n=1}^{k}\frac{(-1)^{n}(\sigma-1)^{n}\gamma_{n}}{n!}+\sum_{n=k+1}^{\infty}\frac{(-1)^{n}(\sigma-1)^{n}\gamma_{n}}{n!}.$ (20)
To prove the non-negativity of (19) we require the last two terms of (20) together to be positive. We have the bounds on $\gamma_{n}$ due to Lavrik [17]:
$|\gamma_{n}|\leq\frac{n!}{2^{n+1}},\quad\quad n=1,2,3,\ldots,$ (21)
hence
$\left|\sum_{n=k+1}^{\infty}\frac{(-1)^{n}(\sigma-1)^{n}\gamma_{n}}{n!}\right|\leq\sum_{n=k+1}^{\infty}\frac{(\sigma-1)^{n}}{2^{n+1}}=-\frac{1}{(\sigma-3)}\left(\frac{\sigma-1}{2}\right)^{k+1}$
by the geometric series with $|1-\sigma|<2$. Taking the lower bound from this, one can then verify that $\phi_{0}(\sigma,k)\geq 0$ for $1<s\leq\sigma_{k}$ by calculating the first $k$ terms. ∎
It is likely that $\phi(\sigma,k)$ is always positive for any $\sigma>1$, although proving that would require a different approach, perhaps one akin to [22, Lemma 5.3]. Our method in Lemmas 10 and 11, while simpler, is restricted by bounds on $|\gamma_{n}|$ which loses some information about sign changes. It is worth noting that bounds sharper than (21) are available, but are more complicated to work with.
Lemma 11.
Let the hypotheses of Lemma 10 be fulfilled. Then for $\sigma$ real, we have
$-\zeta(\sigma)\leq\frac{1}{(\sigma-1)^{2}}+\frac{1}{(\sigma-3)^{2}},\qquad 1<\sigma<3,$
and
$\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq\frac{-\zeta^{\prime}(\sigma)}{\zeta(\sigma)}\leq\frac{\phi_{1}(\sigma,k)}{\sigma-1},\qquad 1<\sigma\leq\sigma_{k},$
where
$\phi_{1}(\sigma,k)=\frac{(\sigma-3)^{2}+(\sigma-1)^{2}}{(\sigma-3)^{2}(1+(\sigma-1)(\phi_{0}(\sigma,k)+\gamma))}.$ (22)
Proof.
The first inequality of the second assertion is trivial. The subsequent inequality is proved by differentiating (20) and repeating a process similar to the proof of Lemma 10 (not splitting the sum for brevity), to obtain
$-\zeta^{\prime}(\sigma)$ $\leq\frac{1}{(\sigma-1)^{2}}+\sum_{n=1}^{\infty}\frac{n(\sigma-1)^{n-1}}{2^{n+1}}$
$=\frac{1}{(\sigma-1)^{2}}+\frac{1}{2(\sigma-1)}\sum_{n=1}^{\infty}n\left(\frac{\sigma-1}{2}\right)^{n}=\frac{1}{(\sigma-1)^{2}}+\frac{1}{(\sigma-3)^{2}}$
for $|1-\sigma|<2$ by identification with the polylogarithm $\textup{Li}_{-1}(z)=z/(1-z)^{2}$. Dividing by the bound in the Lemma 10 proves the result. ∎
These bounds are good for $\sigma$ close to $1$. For instance, we have that
$\phi_{1}(\sigma,3)<1\qquad\text{for}\qquad\sigma\leq 1.83,$ (23)
which is an improvement on Delange’s result [8] for a fixed $\sigma\leq 1.65$. Although the bounds get sharper as $k$ increases, in practice, taking $k\leq 10$ is more than sufficient for small $\sigma$ and any savings gained by subsequent increases in $k$ would be after the fifth decimal place.
Corollary 10.
Let $\phi_{1}(\sigma,k)$ be defined as in (22). For $1<\sigma_{0}\leq\sigma\leq 1.5$, we have
$\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq\frac{-\zeta^{\prime}(\sigma)}{\zeta(\sigma)}\leq\frac{\phi_{2}(\sigma_{0})}{\sigma-1},$
[…]
Proof.
For $k=10$, we have from Lemma 20 that $\phi_{0}(\sigma,10)\geq 0$ in our stated range of $\sigma$. It can also be verified that $\phi_{1}(\sigma,10)$ is decreasing in the region $1<\sigma\leq 1.48$, and therefore $\phi_{1}(\sigma,10)\leq\phi_{1}(\sigma_{0},10)$.
To complete the proof, verify that $\phi_{1}(\sigma,10)\leq 0.851$ in the region $1.48\leq\sigma\leq 1.5$. ∎
Finally, we have a lower bound on $|\zeta(s)|$, for $\sigma>1$, which is better than the trivial bound $|\zeta(s)|\geq\zeta(2\sigma)/\zeta(\sigma)$ over certain ranges of $\sigma$ (for more information, see [19]). This is the $\log t$ analogue of [7, Lemma 8].
Lemma 12.
Let $\kappa,\,\sigma_{1}$ be positive parameters, and $\gamma$ be Euler’s constant. Then for $t\geq t_{0}\geq 3$ and $1+\kappa\leq\sigma_{1}\leq 2$, we have
$\left|\frac{1}{\zeta(1+\kappa+it)}\right|\leq V(\kappa,\sigma_{1},t_{0},\eta)\log t$
with
$V(\kappa,\sigma_{1},t_{0},\eta):=\left(\frac{\exp\left(\gamma\kappa\right)}{\kappa\log t_{0}}\right)^{3/4}\left(1+\frac{\log 2+C(\sigma_{1},t_{0},2,\eta)}{\log t_{0}}\right)^{1/4},$
where $\eta$ and $C(\sigma_{1},t_{0},k,\eta)$ are defined in Corollary 8.
Proof.
Since $\sigma>1$, we use the classical non-negativity argument involving the trigonometric polynomial $2(1+\cos\theta)^{2}=3+4\cos\theta+\cos 2\theta$ (see [26, § 3.3] and [19]). That is, we use $\zeta^{3}(\sigma)|\zeta^{4}(\sigma+it)\zeta(\sigma+2it)|\geq 1$, which implies
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq|\zeta(\sigma)|^{3/4}|\zeta(\sigma+2it)|^{1/4}.$ (24)
Taking $\sigma=1+\kappa$ in (24) gives
$\left|\frac{1}{\zeta(1+\kappa+it)}\right|\leq|\zeta(1+\kappa)|^{3/4}|\zeta(1+\kappa+2it)|^{1/4}.$ (25)
The two factors on the right-hand side can be bounded with Lemmas 9 and 8, respectively, with the aid of the bound
$\log 2t$ $\leq\left(1+\frac{\log 2}{\log t_{0}}\right)\log t.$
This proves the lemma. ∎
4 Estimates for $|\zeta^{\prime}(s)/\zeta(s)|$
4.1 Assuming a partial Riemann hypothesis
In this section, we assume $RH(T)$, and estimate $|\zeta^{\prime}(s)/\zeta(s)|$ in a right-half plane: $\sigma\geq 1/2+\epsilon$, $t_{0}\leq t\leq T$.
We will first prove the result for $T_{0}$ and then make the substitution to $t$ at the end. The idea is to construct discs $|s-s_{0}|\leq 1/2$ which have origins slightly to the right of the $1$-line, centred at $s_{0}=\sigma_{0}+iT_{0}$, where
$\sigma_{0}=1+\epsilon,\qquad e+1/2\leq t_{0}\leq T_{0}\leq T-1/2,$
and $\epsilon\leq 1/2$ is a parameter to be determined later. We then extend their radii into the critical strip, so that Theorem 1 applied to $f(s):=-\zeta^{\prime}(s)/\zeta(s)$ gives us the desired estimates. Due to $RH(T)$, $\zeta(s)$ is analytic and non-zero in the discs.
Choosing a fixed $R^{\prime}<\epsilon$, we have by Corollary 6 that for $1/2+\epsilon\leq\sigma\leq 3/2+\epsilon$,
$\left|{\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(\sigma+iT_{0})\right|\leq M:=\frac{4}{\pi R^{\prime}}\max_{0\leq\theta\leq 2\pi}|\log|\zeta(\sigma+iT_{0}+R^{\prime}e^{i\theta})||,$ (26)
which we estimate, by Lemma 7 with $\delta_{r}=1/2+\epsilon$ and $T_{0}\geq t_{0}\geq e+R^{\prime}$. In the region $1/2\leq\sigma\leq 3/2+2\epsilon$, we have
$|\zeta(s)|\leq A_{0}(T_{0}+R^{\prime})^{1/6}\log(T_{0}+R^{\prime})\leq A_{1}T_{0}^{1/6}\log T_{0},$ (27)
where
$A_{1}$ $=A_{0}\left(1+\frac{R^{\prime}}{t_{0}}\right)^{1/6}\left(1+\frac{\log\left(1+\frac{R^{\prime}}{t_{0}}\right)}{\log t_{0}}\right),$ (28)
with
$A_{0}$ $=0.618\left(1+a_{1}\left(\frac{3}{2}+2\epsilon,1.31,t_{0}-R^{\prime}\right)\right)^{7/6}$ (29)
×$\qquad\qquad\bm{\times}\left(1+a_{0}\left(\frac{3}{2}+2\epsilon,1.31,t_{0}-R^{\prime}\right)\right).$
Combining (27), (29), and (28) with (26) thus gives for $1/2+\epsilon\leq\sigma\leq 3/2+\epsilon$,
$M\leq\frac{4}{\pi R^{\prime}}\left(\frac{1}{6}\log T_{0}+\log\log T_{0}+\log A_{1}\right).$ (30)
Now to bound $|\zeta^{\prime}(s_{0})/\zeta(s_{0})|$, we can use the trivial bound for $\sigma>1$ [12, Lemma 70.1], which gives us
$\left|\frac{\zeta^{\prime}(s_{0})}{\zeta(s_{0})}\right|\leq-\frac{\zeta^{\prime}(\sigma_{0})}{\zeta(\sigma_{0})}<\frac{1}{\sigma_{0}-1}.$ (31)
However, when bounding $|\zeta^{\prime}(s_{0})/\zeta(s_{0})|$ in an interval away from the $1$-line, that is to say, where $\sigma_{0}\leq\sigma\leq 1.5$ with $\sigma_{0}$ not tending to $1$, we can do better than the trivial bound by using Corollary 10. We have for $\sigma_{0}\leq\sigma\leq 1.5$,
$\left|\frac{\zeta^{\prime}(s_{0})}{\zeta(s_{0})}\right|\leq-\frac{\zeta^{\prime}(\sigma_{0})}{\zeta(\sigma_{0})}<\frac{\phi_{2}(\sigma_{0})}{\sigma_{0}-1}.$ (32)
Up to this point, the above with Theorem 1, where $M$ is as in (30), and $R=1/2$, will give a result that holds for
$1+\epsilon-\frac{\alpha}{2}\leq\sigma\leq 1+\epsilon+\frac{\alpha}{2},$
and for a result on any larger $\sigma$, we simply use the trivial bound in (31).
Finally, putting everything together and making the substitution $T_{0}\mapsto t$, we prove our result.
Theorem 2.
Assume $RH(T)$. Let $\alpha$, $\epsilon$, and $R^{\prime}$ be parameters, such that $20<R^{\prime}<\epsilon\leq 1/$ 2 and $10<\alpha<$ 1. In the region
$\sigma\geq 1+\epsilon-\frac{\alpha}{2}\quad\text{and}\quad R^{\prime}+e\leq t_{0}\leq t\leq T-\frac{1}{2},$
we have
$\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq Q_{RH(T)}\log t,$
with
$Q_{RH(T)}$ $=\max\Bigg{\{}\frac{2\alpha}{(1-\alpha)\pi R^{\prime}}\Bigg{(}\frac{1}{6}+B_{t_{0}}+\frac{\log A_{1}}{\log t_{0}}\Bigg{)}$ (33)
$\qquad\qquad+\left(\frac{1+\alpha}{1-\alpha}\right)\frac{\phi_{2}(1+\epsilon)}{\epsilon\log t_{0}},\,\frac{1}{(\epsilon+(\alpha/2))\log t_{0}}\Bigg{\}},$
where $A_{1}$ is defined in (28), $\phi_{2}(\sigma)$ is defined in Corollary 10, and
$B_{t_{0}}=\begin{cases}\frac{1}{e}&\text{ if }t_{0}\leq e^{e},\\ \frac{\log\log t_{0}}{\log t_{0}}&\text{ if }t_{0}>e^{e}.\end{cases}$
Furthermore, by assuming $RH$ instead of $RH(T)$, the upper bound restriction on $t$ can be removed.
Assuming $RH(H_{0})$ allows us to prove Corollary 3, which holds for $t\leq H$.
Proof of Corollary 3.
For $t\geq t_{0}$, in Theorem 2 set $\sigma_{0}=1+\epsilon-\frac{\alpha}{2}$. Then fix $t_{0}$ and optimise (33) over $\epsilon,\,R^{\prime},\,\alpha$, while ensuring that the restrictions set in Theorem 2 and (11) hold. In practice $R^{\prime}$ is taken to be arbitrarily smaller than $\epsilon$, differing only after the sixth decimal place. The final value of $t_{0}=5$ is taken to be the one where (11) holds for all $\sigma_{0}$ in Table 2. This gives the result. ∎
$\sigma_{0}$ $Q_{RH(H_{0})}$ $\epsilon$ $\alpha$
$0.51$ $38889$ $0.004948$ $0.989897$
$0.55$ $1508$ $0.024334$ $0.948668$
$0.6$ $361.726$ $0.047605$ $0.89521$
$0.65$ $153.957$ $0.070599$ $0.841198$
$0.7$ $82.739$ $0.092279$ $0.784558$
$0.75$ $50.441$ $0.113238$ $0.726476$
$0.8$ $33.249$ $0.129407$ $0.658814$
$0.85$ $23.086$ $0.144234$ $0.588468$
$0.9$ $16.601$ $0.160246$ $0.520492$
$0.95$ $12.235$ $0.174136$ $0.448272$
$1$ $9.161$ $0.176533$ $0.353067$
Table 2: Values for $Q_{RH(H_{0})}$ and $\sigma\geq\sigma_{0}$ in Corollary 3, with corresponding $\epsilon,\,\alpha$ (all rounded to six decimal places). Each entry is valid for $5\leq t\leq H$.
4.2 Above the Riemann height
Suppose $t_{0}\geq H$ and let there be a positive real number $d$. As before, we construct discs which have origins slightly to the right of the $1$-line, at $s_{0}=\sigma_{0}+iT_{0}$, where $T_{0}\geq t_{0}$,
$\sigma_{0}=1+\delta_{T_{0}}=1+\frac{d}{\log T_{0}},$ (34)
and extend their radii slightly into the critical strip, and then apply Theorem 1. This time however, we do not have the assumption of $RH(T)$ to help us with potential zeros. We will instead utilise the zero-free region mentioned in Section 2.1 and keep our discs within said region. This requires discs with radii decreasing to 00 at the rate of $1/\log T_{0}$, similar to the zero-free region.
Let
$R_{T_{0}}=2\delta_{T_{0}}=\frac{2d}{\log T_{0}}\qquad\text{and}\qquad r=\alpha R_{T_{0}},\quad 0<\alpha<1.$ (35)
In addition, let $d$, and hence $R_{T_{0}}$, be chosen such that
$\sigma_{0}-R_{T_{0}}=\sigma_{0}-2\delta_{T_{0}}\geq 1-\frac{1}{W_{0}\log(T_{0}+R_{T_{0}})},$
which is fulfilled by
$d\leq\frac{1}{W_{0}}\left(1+\frac{\log\big{(}1+\frac{R_{H}}{H}\big{)}}{\log H}\right)^{-1}.$ (36)
This ensures that the disc $|s-s_{0}|\leq R_{T_{0}}$ lies entirely in the zero-free region, so $f(s)$ is analytic there.
Next, we want to obtain a bound on $M$. This is done by choosing
$R^{\prime}=\frac{1}{2}-\delta_{H}=\frac{1}{2}-\frac{d}{\log H}$
in Lemma 2. Then as long as $d$ is chosen to satisfy (3), i.e.,
$d\leq\frac{1}{W_{0}}\left(1+\frac{\log\big{(}1+\frac{R^{\prime}}{H}\big{)}}{\log H}\right)^{-1},$ (37)
we have for $\sigma\geq 1-\delta_{H}$,
$\left|{\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(\sigma+iT_{0})\right|\leq M:=\frac{4}{\pi R^{\prime}}\max_{0\leq\theta\leq 2\pi}|\log|\zeta(\sigma+iT_{0}+R^{\prime}e^{i\theta})||.$ (38)
Estimating $M$ thus requires a uniform bound of $\zeta(s)$ in the region
$\frac{1}{2}\leq 1-\delta_{T_{0}}-R^{\prime}\leq\sigma\leq 1+3\delta_{T_{0}}+R^{\prime}\leq 1+\frac{1}{2}+2\delta_{H}.$
By Lemma 7 with $\delta_{r}=1/2+2\delta_{H}$, we have for $T_{0}\geq H$, and the region $1/2\leq\sigma\leq 3/2+2\delta_{H}$ that
$|\zeta(s)|\leq A_{2}(T_{0}+R^{\prime})^{1/6}\log(T_{0}+R^{\prime})\leq A_{3}T_{0}^{1/6}\log T_{0},$ (39)
where
$A_{3}$ $=A_{2}\left(1+\frac{R^{\prime}}{H}\right)^{1/6}\left(1+\frac{\log\left(1+\frac{R^{\prime}}{H}\right)}{\log H}\right),$ (40)
with
$A_{2}$ $=0.618\left(1+a_{1}\left(\frac{3}{2}+\frac{2d}{\log H},1.31,H-R^{\prime}\right)\right)^{7/6}$ (41)
×$\qquad\qquad\bm{\times}\left(1+a_{0}\left(\frac{3}{2}+\frac{2d}{\log H},1.31,H-R^{\prime}\right)\right).$
Combining (39), (41), and (40) with (38) thus gives for $1-\delta_{T_{0}}\leq\sigma\leq 1+3\delta_{T_{0}}$,
$M\leq\frac{4}{\pi R^{\prime}}\left(\frac{1}{6}\log T_{0}+\log\log T_{0}+\log A_{3}\right).$ (42)
Now to bound $|\zeta^{\prime}(s_{0})/\zeta(s_{0})|$, we use the trivial bound for $\sigma>1$, given in (31). This yields
$\left|\frac{\zeta^{\prime}(s_{0})}{\zeta(s_{0})}\right|\leq\frac{1}{\sigma_{0}-1}\leq\frac{\log T_{0}}{d}.$ (43)
Up to this point, Theorem 1 will give a result that holds for
$1-\delta_{T_{0}}(2\alpha-1)\leq\sigma\leq 1+\delta_{T_{0}}(2\alpha+1),$ (44)
and for a result on any larger $\sigma$, we simply use (31). At the same time, if we want a result that holds for $\sigma\geq 1-1/W\log T_{0}$, then by the left inequality of (44), we must have
$W=\frac{1}{d(2\alpha-1)}\qquad\text{and}\qquad\alpha=\frac{1}{2}+\frac{1}{2dW}.$ (45)
Finally, putting everything together and making the substitution $T_{0}\mapsto t$, we prove our result.
Theorem 3.
Let $d$ and $W>W_{0}$ be positive parameters, subject to (36), (37), and (45). Let also $10<\alpha<$ 1 be a parameter defined in (45), and depending on $d$ and $W$. In a region
$\sigma\geq 1-\frac{1}{W\log t},\qquad t\geq H$
where $\zeta(\sigma+it)\neq 0$, we have
$\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq Q_{H}\log t,$
where
$Q_{H}=\max\left\{\frac{\lambda}{6}+\frac{1+\alpha}{d(1-\alpha)}+\frac{\lambda(\log\log H+\log A_{3})}{\log H},\,\frac{1}{d(1+2\alpha)}\right\},$
with $A_{3}$ defined in (40),
$\lambda=\frac{8\alpha}{(1-\alpha)\pi R^{\prime}},\qquad\text{and}\qquad R^{\prime}=\frac{1}{2}-\frac{d}{\log H}.$
4.3 Combining both cases $t\leq H$ and $t\geq H$
Proof of Corollary 1.
For the case of $t\geq H$, with a fixed choice of $W$ as defined in (45), it remains to optimise $Q_{H}$ in Theorem 3 over $\alpha$ and $d$, subject to the conditions mentioned in said theorem, while also checking that (11) holds. We found that the optimal value for $d$ in each case was approximately $d=0.179898\ldots$, and any differences in the final result caused by varying $d$ further occurred only after the sixth decimal place. We note also that the optimal $d$ is one that is as close to $1/W_{0}$ as possible while still satisfying (36) and (37).
On the other hand, a simple calculation shows that
$0.821-\frac{1}{W_{0}\log t}>$ 0.82
for $t\geq t_{0}=e$. Now, for $\sigma\geq 0.8$ and $5\leq t\leq H$, Corollary 3 and Table 2 asserts that
$\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq Q_{RH(H_{0})}\log t=33.249\log t,$
which is smaller than any value of optimised $Q_{H}$ found. Thus this same value of $Q_{H}$ holds for all $t\geq 5$. ∎
We also prove Corollary 2, especially relevant for $\sigma=1$.
Proof of Corollary 2.
First note that Corollary 3 and Table 2 has for $\sigma\geq 1$ and $5\leq t\leq H$,
$\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq Q_{RH(H_{0})}\log t=9.161\log t.$
Then for $t\geq H$, the proof is identical to that of Theorem 3 and Corollary 1, except we can restrict the size of $\alpha$ even further, since we now require $\sigma_{0}-\alpha R_{T_{0}}\geq 1$, which is equivalent to $\alpha\leq 1/2$. In order for the result to hold for all $\sigma\geq 1$, we choose $\alpha=1/2$, so in this case,
$Q_{H}=\max\left\{\frac{8}{6\pi R^{\prime}}+\frac{3}{d}+\frac{8(\log\log H+\log A_{3})}{\pi R^{\prime}\log H},\,\frac{1}{2d}\right\}.$
As before, we then optimise over $d$, while ensuring that we fulfil the conditions in Theorem 3 and (11). The optimal value found was $d=0.179898$. ∎
$W$ $Q$ $\alpha$ $W$ $Q$ $\alpha$
$5.559$ $450041$ $0.999972$ $5.59$ $4458$ $0.997200$
$5.56$ $106223$ $0.999882$ $5.6$ $3383$ $0.996312$
$5.5601$ $98674$ $0.999873$ $5.7$ $1002$ $0.987604$
$5.5602$ $92137$ $0.999864$ $5.8$ $594.087$ $0.979198$
$5.5603$ $86414$ $0.999855$ $5.9$ $425.318$ $0.971076$
$5.5604$ $81354$ $0.999846$ $6$ $333.034$ $0.963224$
$5.5605$ $76859$ $0.999837$ $6.1$ $274.847$ $0.955630$
$5.561$ $60223$ $0.999792$ $6.2$ $234.806$ $0.948282$
$5.562$ $42026$ $0.999703$ $6.3$ $205.568$ $0.941166$
$5.563$ $32278$ $0.999613$ $6.4$ $183.281$ $0.934273$
$5.564$ $26201$ $0.999523$ $6.5$ $165.729$ $0.927592$
$5.565$ $22051$ $0.999433$ $6.75$ $134.741$ $0.911755$
$5.566$ $19036$ $0.999343$ $7$ $114.502$ $0.897049$
$5.567$ $16748$ $0.999254$ $8$ $74.999$ $0.847418$
$5.568$ $14951$ $0.999164$ $9$ $58.454$ $0.808816$
$5.569$ $13502$ $0.999074$ $10$ $49.359$ $0.777935$
$5.57$ $12310$ $0.998985$ $11$ $43.607$ $0.752668$
$5.575$ $8541$ $0.998537$ $12$ $39.642$ $0.731612$
$5.58$ $6542$ $0.998091$ $13$ $36.742$ $0.713796$
Table 3: Values for $Q$ in Corollary 1, with corresponding $W$ and $\alpha$ (rounded to six decimal places). Each entry is valid for $t\geq 5$.
5 Estimates for $|1/\zeta(s)|$
Moving from a bound on the logarithmic derivative of $\zeta$ to one on the reciprocal of $\zeta$ is done in the usual way (see [26, Theorem 3.11]). It revolves around estimating the right-hand side of
$\log\left|\frac{1}{\zeta(\sigma+it)}\right|=-\textup{Re}\log\zeta\left(1+\delta_{1}+it\right)+\int_{\sigma}^{1+\delta_{1}}\textup{Re}\left(\frac{\zeta^{\prime}}{\zeta}(x+it)\right)\text{d}x$ (46)
for some $\delta_{1}>0$.
The estimation of the first term on the right-hand side can be improved using the trigonometric polynomial (see [3, Proposition A.2], [19]), while in the past the above integral has been estimated by bounding the real part of the logarithmic derivative by a uniform bound $|\zeta^{\prime}(s)/\zeta(s)|$. Here some savings can be made by splitting up the interval $[\sigma,1+\delta_{1}]$ and using the sharpest bounds for $|\zeta^{\prime}(s)/\zeta(s)|$ in the appropriate range, rather than one uniform bound for the entire interval (see [7, Lemma 10]).
However, we show that by using directly a bound on the real part instead, through Lemma 2, we are able to do much better. Furthermore, through a trick utilising the trigonometric polynomial, we are able to get an slight asymptotic improvement on the bounds, at the cost of a slightly worse constant. As before, the proofs will be split into two cases, $t\leq H$ and $t\geq H$.
5.1 Assuming a partial Riemann hypothesis
Theorem 4.
Assume $RH(T)$. Let $\gamma$ be Euler’s constant and let there be positive parameters $d_{1}$, $\sigma_{1}$, $W$, $d$, such that $W=1/d$, and $1+d_{1}/\log t_{0}\leq\sigma_{1}\leq 2$. Suppose also $R^{\prime}=1/2-d/\log t_{0}$. Then in the region
$\sigma\geq 1-\frac{1}{W\log t}\qquad\text{and}\qquad R^{\prime}+e\leq t_{0}\leq t\leq T-\frac{1}{2},$
we have
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq Y_{RH(T)}(d,d_{1},\sigma_{1},t_{0},\eta)\log t,$
where
$Y_{RH(T)}$ $(d,d_{1},\sigma_{1},t_{0},\eta)=\max\Bigg{\{}\frac{1}{d_{1}}+\frac{1}{\log t_{0}},$
$\exp\Bigg{(}\frac{4(d_{1}+d)}{\pi R^{\prime}}\left(\frac{1}{6}+B_{t_{0}}+\frac{\log A_{4}}{\log t_{0}}\right)\Bigg{)}V\left(\frac{d_{1}}{\log t_{0}},\sigma_{1},t_{0},\eta\right)\Bigg{\}},$
with
$A_{4}$ $=\left(1+\frac{R^{\prime}}{t_{0}}\right)^{1/6}\left(1+\frac{\log\left(1+\frac{R^{\prime}}{t_{0}}\right)}{\log t_{0}}\right)\left(1+a_{0}\left(\frac{3}{2}+\frac{d_{1}-d}{\log t_{0}},1.31,t_{0}-R^{\prime}\right)\right)$
×$\qquad\qquad\bm{\times}0.618\left(1+a_{1}\left(\frac{3}{2}+\frac{d_{1}-d}{\log t_{0}},1.31,t_{0}-R^{\prime}\right)\right)^{7/6},$
$B_{t_{0}}$ defined in Theorem 2, and $V(d_{1}/\log t_{0},\sigma_{1},t_{0},\eta),\,\eta$ defined in Lemma 12.
Furthermore, by assuming $RH$ instead of $RH(T)$, the upper bound restriction on $t$ can be removed.
Proof.
Let $\delta_{t}:=d/\log t$ and $\delta_{1,t}:=d_{1}/\log t$, and first suppose $1-\delta_{t}\leq\sigma\leq 1+\delta_{1,t}$. Since by $RH(T)$ there are no zeros for $\sigma\geq 1/2$, we do not have to check the condition in (3). The same argument in Section 4.1 with Lemma 2 leading to (30), requires a uniform bound on $|\zeta(s)|$ in the region
$\frac{1}{2}\leq 1-\delta_{t}-R^{\prime}\leq\sigma\leq 1+\delta_{1,t}+R^{\prime}\leq\frac{3}{2}+\frac{d_{1}-d}{\log t_{0}},$
and in return gives us
$\left|{\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(\sigma+it)\right|\leq\frac{4}{\pi R^{\prime}}\left(\frac{1}{6}+B_{t_{0}}+\frac{\log A_{4}}{\log t_{0}}\right)\log t$
uniformly for $1-\delta_{t}\leq\sigma\leq 1+\delta_{1,t}$. Thus we have from (46) that
$\log\left|\frac{1}{\zeta(\sigma+it)}\right|\leq\log\left|\frac{1}{\zeta(1+\delta_{1,t}+it)}\right|+\int_{1-\delta_{t}}^{1+\delta_{1,t}}\left|{\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(x+it)\right|\text{d}x,$ (47)
where the integral is bounded from above by
$\frac{4(d_{1}+d)}{\pi R^{\prime}}\left(\frac{1}{6}+B_{t_{0}}+\frac{\log A_{4}}{\log t_{0}}\right).$
To bound the first term on the right-hand side of (47), we use Lemma 12 with $\kappa=\delta_{1,t}=d_{1}/\log t$ to see that
$\left|\frac{1}{\zeta(1+\delta_{1,t}+it)}\right|\leq V\left(\frac{d_{1}}{\log t_{0}},\sigma_{1},t_{0},\eta\right)\log t,$ (48)
Hence (47) becomes
$\left|\frac{1}{\zeta(\sigma+it)}\right|$ $\leq\exp\Bigg{(}\frac{4(d_{1}+d)}{\pi R^{\prime}}\left(\frac{1}{6}+B_{t_{0}}+\frac{\log A_{4}}{\log t_{0}}\right)\Bigg{)}V\left(\frac{d_{1}}{\log t_{0}},\sigma_{1},t_{0},\eta\right)\log t,$
for $t_{0}\leq t\leq T-1/2$ and $1-\delta_{t}\leq\sigma\leq 1+\delta_{1,t}$.
Now we turn to the case $\sigma\geq 1+\delta_{1,t}$. Here we use the trivial bound
$|\zeta(s)|\leq\zeta(\sigma)\leq\frac{\sigma}{\sigma-1},\qquad(\sigma>1),$ (49)
where $\sigma/(\sigma-1)$ is decreasing in $\sigma$. Together with the trigonometric inequality from (25), we see that
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq\zeta(1+\delta_{1,t})\leq\left(\frac{1}{d_{1}}+\frac{1}{\log t_{0}}\right)\log t.$ (50)
We reiterate that although there are sharper bounds available to use here, this is sufficient for our purposes as the case where $\sigma\leq 1+\delta_{1,t}$ is dominant when computing constants.
Note that the results proved hold also if one assumes $RH$, and allow us to remove the upper bound restriction on $t$, as explained at the end of the proof of Theorem 2. ∎
5.2 Above the Riemann height
Now we consider $t\geq H$ and prove the following.
Theorem 5.
Let $\gamma$ be Euler’s constant and let there be positive parameters $d_{1}$, $\sigma_{1}$, $W$, $d$, such that $W=1/d$, and $1+d_{1}/\log H\leq\sigma_{1}\leq 2$. Suppose also $R^{\prime}=1/2-d/\log H$, and $W>W_{0}$ is chosen such that $d$ satisfies (37). Then in the region
$\sigma\geq 1-\frac{1}{W\log t},\qquad t\geq H,$
we have
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq Y_{H}(d,d_{1},\sigma_{1},\eta)\log t,$
where
$Y_{H}$ $(d,d_{1},\sigma_{1},\eta)=\max\Bigg{\{}\frac{1}{d_{1}}+\frac{1}{\log H},$
$\exp\Bigg{(}\frac{4(d_{1}+d)}{\pi R^{\prime}}\left(\frac{1}{6}+\frac{\log\log H}{\log H}+\frac{\log A_{5}}{\log H}\right)\Bigg{)}V\left(\frac{d_{1}}{\log H},\sigma_{1},H,\eta\right)\Bigg{\}},$
with
$A_{5}$ $=\left(1+\frac{R^{\prime}}{H}\right)^{1/6}\left(1+\frac{\log\left(1+\frac{R^{\prime}}{H}\right)}{\log H}\right)\left(1+a_{0}\left(\frac{3}{2}+\frac{d_{1}-d}{\log H},1.31,H-R^{\prime}\right)\right)$
×$\qquad\qquad\bm{\times}0.618\left(1+a_{1}\left(\frac{3}{2}+\frac{d_{1}-d}{\log H},1.31,H-R^{\prime}\right)\right)^{7/6},$
and $V(d_{1}/\log H,\sigma_{1},H,\eta),\,\eta$ are defined in Lemma 12. Admissible values of $Y_{H}$ with corresponding $W$ are presented in Table 4.
Proof.
The proof is analogous to that of Theorem 4. Let $\delta_{t}:=d/\log t$ and $\delta_{1,t}:=d_{1}/\log t$, and first suppose $1-\delta_{t}\leq\sigma\leq 1+\delta_{1,t}$. As we do not assume $RH(T)$, we have to check that (3) holds, and we do this by our hypothesis of $d$ satisfying (37). The same argument in Section 4.2 with Lemma 2 leading to (42), requires a uniform bound on $|\zeta(s)|$ in the region
$\frac{1}{2}\leq 1-\delta_{t}-R^{\prime}\leq\sigma\leq 1+\delta_{1,t}+R^{\prime}\leq\frac{3}{2}+\frac{d_{1}-d}{\log H},$
and in return gives us
$\left|{\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(\sigma+it)\right|\leq\frac{4}{\pi R^{\prime}}\left(\frac{1}{6}+\frac{\log\log H}{\log H}+\frac{\log A_{5}}{\log H}\right)\log t$
uniformly for $1-\delta_{t}\leq\sigma\leq 1+\delta_{1,t}$.
The rest of the proof then follows that of Theorem 4, and it remains to optimise $Y_{H}$ for a fixed choice of $W=1/d$, over $d_{1}$, $\sigma_{1}$ and $\eta$, subject to the conditions mentioned in the statement of the theorem, while also checking that (11) holds. ∎
$W$ $Y_{H}$ $d_{1}$ $\sigma_{1}$ $\eta$
$5.5587$ $2.307$ $1.059544$ $1.412209$ $3.464102$
$6$ $2.284$ $1.060503$ $1.608124$ $3.775135$
$7$ $2.244$ $1.062065$ $1.604395$ $3.772048$
$8$ $2.214$ $1.063547$ $1.895304$ $3.621918$
$9$ $2.191$ $1.064570$ $1.399356$ $3.464227$
$10$ $2.173$ $1.065554$ $1.135564$ $3.65928$
$11$ $2.158$ $1.065999$ $1.43905$ $3.446264$
$12$ $2.146$ $1.066605$ $1.345521$ $3.769427$
$13$ $2.136$ $1.067103$ $1.345557$ $3.77319$
Table 4: Values for $Y_{H}$ in Theorem 5, with corresponding $W$, and parameters (rounded to six decimal places) $d_{1}$, $\sigma_{1}$, and $\eta$. Each entry is valid for $t\geq H$.
5.3 Combining both cases $t\leq H$ and $t\geq H$
Proof of Corollary 4.
For the first assertion of the corollary, we begin by assuming $RH(H_{0})$ and dealing with the case $t_{0}\leq t\leq H$.
To obtain $Y(W,5)$, fix $t_{0}$ and $W=1/d$ in Theorem 4 and optimise $Y_{RH(H_{0})}$ over $d_{1}$, $\sigma_{1}$, and $\eta$. At the same time ensure that the restrictions set in Theorem 4 and Corollary 8 hold, and raise $t_{0}$ if (11) is not satisfied. We find that the smallest value of $t_{0}$ such that (11) is fulfilled for all entries in Tables 6 and 5 is $t_{0}=5$. Then we compare the resulting values with the corresponding $Y_{H}$ (for $t\geq H$) in Theorem 5 and Table 4, taking the larger of the two.
[…]
For the second assertion, i.e., when $\sigma=1$, we set $d=0$ and repeat the above process using Theorems 4 (under $RH(H_{0})$) and 5 (unconditionally). This allows us to obtain a result for $\sigma\geq 1$. We have
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq 2.018\log t\quad\text{for}\quad t\geq H,$ (51)
[…]
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq 2.751\log t\quad\text{for}\quad 2\exp(e^{2})\leq t\leq H,$ (52)
[…]
Finally, combine (51) and (52) with the following two bounds. First, we utilise Lemma 7, which was proven using numerical methods. Taking $\sigma=1$ in said lemma, we derive
$\left|\frac{1}{\zeta(1+it)}\right|\leq\frac{8}{\log(500)}\log t\leq 1.288\log t\quad\text{for}\quad 500\leq t\leq 2\exp(e^{2}).$
Second, a computation from the proof of [3, Proposition A.2], implemented in interval arithmetic, states
$\left|\frac{1}{\zeta(1+it)}\right|\leq 2.079\log t\quad\text{for}\quad 2\leq t\leq 500.$
This completes the proof. ∎
When $\sigma$ is fixed (in this case $\sigma=1$), the limiting factor for the constant in the second assertion of Corollary 4 comes from the case $2\exp(e^{2})\leq t\leq H$. In particular, the lower threshold of $2\exp(e^{2})$ is what determines the constant $2.751$. Using numerical methods, it is possible to push this threshold higher and hence improve (52). To illustrate this idea, following the proof of Corollary 4, we can prove unconditionally that for $t\geq 10^{6}$,
$\left|\frac{1}{\zeta(1+it)}\right|\leq 2.391\log t,$ (53)
which is an improvement over what one gets from [5, Theorem 1.1].
We make the remark that we include our computed values of $\sigma_{1}$ and $\eta$ in Tables 4, 5, and 6, for completeness sake. In practice, as we limit the overall constant to three decimal places, $d_{1}$ is the main parameter we optimise over. For $\sigma_{1}$ and $\eta$, any sufficiently large values will do, and further optimisations here will only improve the final constant from the fourth decimal place onwards.
$t_{0}$ $Y(5.5587,t_{0})$ $d_{1}$ $\sigma_{1}$ $\eta$
$10$ $8.101$ $0.307606$ $1.330692$ $3.068541$
$10^{2}$ $4.339$ $0.542825$ $1.851563$ $3.417699$
$10^{3}$ $3.632$ $0.648814$ $1.282661$ $3.465621$
$10^{4}$ $3.264$ $0.725026$ $1.748467$ $3.463629$
$10^{5}$ $3.021$ $0.787049$ $1.208003$ $3.464054$
$10^{7}$ $2.711$ $0.885011$ $1.101137$ $3.464109$
$10^{9}$ $2.518$ $0.960512$ $1.272668$ $3.445896$
$10^{11}$ $2.383$ $1.021396$ $1.655004$ $3.464097$
Table 5: Values for $Y(5.5587,t_{0})$ in Corollary 4, with corresponding $t_{0}$, and parameters for the case $t_{0}\leq t\leq H$ (rounded to six decimal places): $d_{1}$, $\sigma_{1}$, and $\eta$. Each entry is valid for $t\geq t_{0}$ and $W=5.5587$.
$W$ $Y(W,5)$ $d_{1}$ $\sigma_{1}$ $\eta$
$5.5587$ $18.060$ $0.170031$ $1.105647$ $2.784753$
$6$ $16.722$ $0.173040$ $1.107519$ $2.781151$
$7$ $14.654$ $0.178370$ $1.110828$ $2.783844$
$8$ $13.339$ $0.182334$ $1.113292$ $2.786533$
$9$ $12.433$ $0.185403$ $1.115198$ $2.782321$
$10$ $11.771$ $0.187839$ $1.116712$ $2.782458$
$11$ $11.268$ $0.189804$ $1.117933$ $2.786745$
$12$ $10.872$ $0.191448$ $1.118954$ $2.786457$
$13$ $10.553$ $0.192838$ $1.119819$ $2.786611$
Table 6: Values for $Y(W,5)$ in Corollary 4, with corresponding $W$, and parameters for the case $t_{0}\leq t\leq H$ (rounded to six decimal places): $d_{1}$, $\sigma_{1}$, and $\eta$. Each entry is valid for $t\geq t_{0}=5$.
5.4 An asymptotic improvement
In this section we prove an asymptotic improvement over the bounds given earlier in Section 5, at the expense of a worse constant. For the sake of clarity, we first outline the general idea here, before proving the next two theorems.
Assume we have the bound
$\max_{1-\delta\leq\sigma\leq 1+\delta_{1}}\left|{\textup{Re}}\left(\frac{\zeta^{\prime}}{\zeta}(\sigma+it)\right)\right|\leq C\log t,$
for $t\geq t_{0}$ and $1-\delta\leq\sigma\leq 1+\delta_{1}$, where $\delta=d/\log t$ and $\delta_{1}=d_{1}/\log t$, as before. Then from (46), we have
$\left|\frac{1}{\zeta(\sigma+it)}\right|$ $\leq\left|\frac{1}{\zeta(1+\delta_{1}+it)}\right|\exp\left((\delta_{1}+\delta)\max_{1-\delta\leq\sigma\leq 1+\delta_{1}}\left|{\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(\sigma+it)\right|\right)$
$\leq\left|\frac{1}{\zeta(1+\delta_{1}+it)}\right|\exp\left((d_{1}+d)C\right).$ (54)
Notice that since the exponential factor is $O(1)$, the overall order is determined by the first factor.
With the same argument involving the trigonometric inequality (24),
$\left|\frac{1}{\zeta(1+\delta_{1}+it)}\right|\leq|\zeta(1+\delta_{1})|^{3/4}|\zeta(1+\delta_{1}+2it)|^{1/4},$ (55)
and at the expense of a worse constant, we can estimate the second factor in (55) differently from the previous sections. We do this by using Lemma 6, which states that $|\zeta(s)|\leq C_{3}(\log t)^{2/3}$ for $\sigma\geq 1$. Along with Lemma 9, this leads to
$\left|\frac{1}{\zeta(1+\delta_{1}+it)}\right|\leq(C_{4}\log t)^{3/4}(C_{3}(\log t)^{2/3})^{1/4},$
which applied to (5.4) ultimately furnishes the bound
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq C_{5}(\log t)^{11/12}.$
Theorem 6.
Assume $RH(T)$. Let $\gamma$ be Euler’s constant and let there be positive parameters $d$, $d_{1}$, $W$, such that $W=1/d$, and $d_{1}/\log t_{0}\leq 1$. Suppose also $R^{\prime}=1/2-d/\log t_{0}$. Then in the region
$\sigma\geq 1-\frac{1}{W\log t}\qquad\text{and}\qquad R^{\prime}+e\leq t_{0}\leq t\leq T-\frac{1}{2},$
we have
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq Y^{\prime}_{RH(T)}(d,d_{1},t_{0})(\log t)^{11/12},$
where
$Y^{\prime}_{RH(T)}$ $(d,d_{1},t_{0})=C_{3}^{1/4}\cdot\max\Bigg{\{}\left(\frac{1}{d_{1}}+\frac{1}{\log t_{0}}\right)^{3/4},$
$\exp\Bigg{(}\frac{4(d_{1}+d)}{\pi R^{\prime}}\left(\frac{1}{6}+B_{t_{0}}+\frac{\log A_{4}}{\log t_{0}}\right)\Bigg{)}\left(\frac{e^{\gamma d_{1}/\log t_{0}}}{d_{1}}\right)^{3/4}\Bigg{\}},$
with $A_{4}$ defined in Theorem 4, $B_{t_{0}}$ defined in Theorem 2, and $C_{3}$ defined in Lemma 6.
Furthermore, by assuming $RH$ instead of $RH(T)$, the upper bound restriction on $t$ can be removed.
Proof.
The proof is identical to that of Theorem 4, except that for $\sigma\leq 1+\delta_{1,t}$, by Lemmas 6 and 9, we can replace (48) with
$\left|\frac{1}{\zeta(1+\delta_{1,t}+it)}\right|$ $\leq|\zeta(1+\delta_{1,t})|^{3/4}|\zeta(1+\delta_{1,t}+2it)|^{1/4}$
$\leq\left(\frac{e^{\gamma(d_{1}/\log t_{0})}}{d_{1}}\log t\right)^{3/4}\left(C_{3}(\log t)^{2/3}\right)^{1/4}.$
Meanwhile, for $\sigma\geq 1+\delta_{1,t}$, the same argument with (24) and (49) leads to
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq\left(\left(\frac{1}{d_{1}}+\frac{1}{\log t_{0}}\right)\log t\right)^{3/4}\left(C_{3}(\log t)^{2/3}\right)^{1/4}$
in place of (50). ∎
Theorem 7.
Let $\gamma$ be Euler’s constant and let there be positive parameters $d$, $d_{1}$, $W$, such that $W=1/d$, and $d_{1}/\log H\leq 1$. Suppose also $R^{\prime}=1/2-d/\log H$, and $W>W_{0}$ is chosen such that $d$ satisfies (37). Then in the region
$\sigma\geq 1-\frac{1}{W\log t},\qquad t\geq H,$
we have
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq Y^{\prime}_{H}(d,d_{1})(\log t)^{11/12},$
where
$Y^{\prime}_{H}$ $(d,d_{1})=C_{3}^{1/4}\cdot\max\Bigg{\{}\left(\frac{1}{d_{1}}+\frac{1}{\log H}\right)^{3/4},$
$\exp\Bigg{(}\frac{4(d_{1}+d)}{\pi R^{\prime}}\left(\frac{1}{6}+\frac{\log\log H}{\log H}+\frac{\log A_{5}}{\log H}\right)\Bigg{)}\left(\frac{e^{\gamma d_{1}/\log H}}{d_{1}}\right)^{3/4}\Bigg{\}},$
with $A_{5}$ defined in Theorem 5, and $C_{3}$ defined in Lemma 6.
Proof.
The proof is similar to that of Theorem 6, except we do not assume $RH(T)$. Then the process follows the proof of Theorem 5, making similar changes as with the earlier case $t\leq H$ in Theorem 6. ∎
Combining Theorems 6 and 7, we prove Corollary 5, a result which holds for all $t\geq t_{0}\geq 5$.
Proof of Corollary 5.
The proof is analogous to that of Corollary 4, where the final result is obtained by optimising over the appropriate parameters.
For the result on $\sigma=1$, the limiting region is again $2\exp(e^{2})\leq t\leq H$, where we have
$\left|\frac{1}{\zeta(1+it)}\right|\leq 7.475(\log t)^{11/12},$
with $d_{1}=0.72042$. ∎
$t_{0}$ $Y^{\prime}(5.5587,t_{0})$ $d_{1}$ $W$ $Y^{\prime}(W,5)$ $d_{1}$
$10$ $21.123$ $0.307588$ $5.5587$ $47.054$ $0.170052$
$10^{2}$ $11.643$ $0.542830$ $6$ $43.568$ $0.173052$
$10^{3}$ $9.842$ $0.648792$ $7$ $38.179$ $0.178401$
$10^{4}$ $8.885$ $0.725028$ $8$ $34.753$ $0.182368$
$10^{5}$ $8.247$ $0.787048$ $9$ $32.391$ $0.185428$
$10^{7}$ $7.425$ $0.885011$ $10$ $30.668$ $0.187861$
$10^{9}$ $6.907$ $0.960531$ $11$ $29.356$ $0.18984$
$10^{11}$ $6.545$ $1.021397$ $12$ $28.325$ $0.191481$
$H$ $6.340$ $1.059544$ $13$ $27.494$ $0.192568$
Table 7: Values for $Y^{\prime}(W,t_{0})$ in Corollary 5, with corresponding $W$ or $t_{0}$, and parameter $d_{1}$ for the case $t_{0}\leq t\leq H$ (rounded to six decimal places). Each entry is valid for $t\geq t_{0}$.
6 Concluding Remarks
Here we state some possible ways to improve our estimates. The most impactful would be having a larger zero-free region or a higher Riemann height. Using an asymptotically stronger zero-free region is also helpful. See for instance [7], where a Littlewood-style zero-free region is used to obtain bounds of order $\log t/\log\log t$. Minor improvements can be made by a more careful treatment of the sum over zeros and the integral arising from the application of Lemma 1. Other possibilities are: an improvement on the Borel–Carathéodory by specialising it more to this problem, proving a uniform (in $\sigma$) bound for $\arg\zeta(s)$, or using stronger estimates for $\zeta(s)$ under $RH$ or $RH(T)$.
It is evident from the results that the closer $s$ is to the boundary of the zero-free region, and hence a potential zero, the larger the constants. However, the large constants can also be attributed in part to the factor $(R+r)/(R-r)$ in Theorem 1, which increases rapidly as $r$ approaches $R$. This contributes significantly to the overall estimate, even if $|f(s_{0})|$ is small, and seems to suggest that Theorem 1 in its current form might not be optimal for this problem.
The main advantage of our method using Lemma 1 is that it eliminates the need for Theorem 1 in estimating $|1/\zeta(s)|$, and also reduces the number of applications of the theorem when estimating $|\zeta^{\prime}(s)/\zeta(s)|$. It might be possible to remove this entirely, by directly bounding the real and imaginary part of $|\zeta^{\prime}(s)/\zeta(s)|$, as one can easily prove an analogue of Lemma 1 which relates $\textup{Im}(f^{\prime}/f)$ to $\arg f$. To use such a result in the case of $f(s)=\zeta(s)$, besides dealing with the corresponding sum over zeros, a uniform bound (over a range of $\sigma$) on $S_{\sigma}(t):=(1/\pi)\arg\zeta(\sigma+it)$ is needed.
To get unconditional results we have (due to knowledge of the Riemann height) assumed a partial Riemann hypothesis to prove some of our results. As an example, for $W=5.5587$, the condition (37) in Theorem 5 is only fulfilled for $t\geq 3.22\cdot 10^{9}$. The conditions on $d$ (which depends on $t$) are what prevents us from taking $W$ too small, and this is where the Riemann height and $RH(T)$ helps us. In fact we only need a ‘weak’ version of $RH$, for instance in (26), that up to height $T$, no zeros lie to the right of $1/2-R^{\prime}$. The downside to this is that we have not used better available bounds that require the full strength of $RH$. For example, under $RH$ we could have used that $\zeta(1/2+it)=O(\exp(C\log t/\log\log t))$ for some $C>0$ to prove something stronger (see [26, §14], [24], [23], and [6] for more details).
Even in the unconditional results used, there is room for improvement. As an illustration, Corollary 5, though asymptotically superior, beats Corollary 4 only when $t$ is large. In the case of $\sigma=1$, this happens for $t\geq\exp(161975)$. This is in part due to the relatively large constants in Corollary 5, and a direct way of lowering these constants is an improvement in the bound $|\zeta(1+it)|$, which correspondingly lowers $C_{3}$ in Lemma 6.
Finally, we conclude with a discussion on the trigonometric inequality (24) used in many of our above proofs, in particular when bounding $|1/\zeta(s)|$ for $\sigma>1$. Under certain conditions, this method is superior to using the trivial bound. In general, when studying zero-free regions for $\zeta(s)$, increasing the degree of the non-negative trigonometric polynomial improves the results one gets. Surprisingly, the same cannot be said in our setting (with the exception of results in Section 5.4), and the classical polynomial we used is actually optimal. Take for instance Lemma 12, which is used in Theorems 4 and 5. Here the overall order of the bound is $O(\log t)$, and in this scenario the article [19] has an in-depth discussion on why the classical polynomial is best.
On the other hand, in Theorems 6 and 7, the question is whether higher degree polynomials lead to a reduction in the exponent $11/12$, even if overall constants get worse. This is an interesting question to consider, since in this case the answer is less obvious, although preliminary calculations suggest that the classical polynomial is still the optimal one.
Acknowledgements
We thank Tim Trudgian for his guidance throughout the writing of this paper. We are also very grateful to Bryce Kerr, Olivier Ramaré, and Aleksander Simonič for the helpful discussions.