### v4

# Explicit estimates for the logarithmic derivative and the reciprocal of the Riemann zeta function

Abstract
In this article, we give explicit bounds of order $\log t$ for $\sigma$ close to $1$, for two quantities: $|\zeta^{\prime}(\sigma+it)/\zeta(\sigma+it)|$ and $|1/\zeta(\sigma+it)|$. We correct an error in the literature, and especially in the case of $|1/\zeta(\sigma+it)|$, also provide improvements in the constants. Using an argument involving the trigonometric polynomial, we additionally provide a slight asymptotic improvement within the classical zero-free region: $1/\zeta(\sigma+it)\ll(\log t)^{11/12}$. The same method applied to the Korobov–Vinogradov zero-free region gives a new record: the unconditional bound $1/\zeta(\sigma+it)\ll(\log t)^{2/3}(\log\log t)^{1/4}$.
1 Introduction
In the study of the Riemann zeta function $\zeta(s)$ and related functions, one often requires upper bounds for the logarithmic derivative of $\zeta(s)$ and the reciprocal of $\zeta(s)$. For example, these come up in the Perron formula in the process of estimating the Chebyshev function $\psi(x)=\sum_{n\leq x}\Lambda(n)$, where $\Lambda(n)$ is the von Mangoldt function, or $M(x)=\sum_{n\leq x}\mu(n)$, where $\mu(n)$ is the Möbius function. The reader is referred to [29], [4], [9], [26, (3.14)] for more information on this.
The chief purpose of this article is to prove explicit upper bounds for $\zeta^{\prime}(s)/\zeta(s)$ and $1/\zeta(s)$. Along the way, we also provide in §3 a collection of other improved bounds on quantities involving $\zeta(s)$, which might be useful for anyone seeking to do explicit computations in the area.
Our main results will be presented in detail in §2, but let us first give a brief history of results on the two quantities $\zeta^{\prime}(s)/\zeta(s)$ and $1/\zeta(s)$. Trudgian in [29] gives an explicit version of all the lemmas in [26, §3.9], which were originally due to Landau. This then allowed Trudgian to obtain explicit bounds on $\zeta^{\prime}(s)/\zeta(s)$ and $1/\zeta(s)$ of order $O(\log t)$. This was then extended in [15] and [7] to explicit bounds of order $O(\log t/\log\log t)$. However, estimates for the two quantities mentioned above tend to have large constants, often rendering them unsuitable for practical use. It is a non-trivial task to reduce these constants. This is due to the fact that the factor of $\zeta(s)$ in the denominator means that such estimates can only be valid in a zero-free region, and the constants rapidly increase in size as we approach the boundary of the region, i.e., $\zeta(s)\to 0$.
While the bounds on $|\zeta^{\prime}(s)/\zeta(s)|$ given in this paper might seem worse compared to [29], we take the opportunity to point out an error in that paper. In the quantity $A$ of [29, Lemma 4], there should be a factor of $1/c$ instead of $1/2c$. This leads to the overall constants being better than they should be. For instance, in Table $2$ of said paper, which shows bounds of the form $|\zeta^{\prime}(s)/\zeta(s)|\leq R_{1}\log t$ for $\sigma\geq 1-1/W\log t$, one should have $(W,R_{1})=(6,538.16)$ and $(12,59.49)$, instead of $(6,382.58)$ and $(12,40.14)$, respectively. This also impacts the subsequent bounds on $|1/\zeta(s)|$ ($R_{2}$ in [29]), which are derived from $R_{1}$. In this article, we have corrected the erroneous bounds and even slightly improved upon the (corrected) bounds.
Furthermore, by adapting a method of [7, §5] and being more careful in our derivation on bounds of $|1/\zeta(s)|$, we obtain improvements in some (erroneously better than they should be) bounds in [29]. In addition, when $\sigma=1$, the final assertion of Corollary 4 is significantly better than [3, Proposition A.2].
Finally, a novel argument utilising a combination of the classical non-negative trigonometric polynomial and a uniform bound of the form $\zeta(s)\ll(\log t)^{2/3}$ for $\sigma\geq 1$, allows us to get an asymptotic improvement in the case of $1/\zeta(s)$. In particular, working within zero-free regions of the forms
$\sigma\geq 1-\frac{1}{Z_{1}\log t},\qquad\sigma\geq 1-\frac{\log\log t}{Z_{2}\log t},\qquad\sigma\geq 1-\frac{1}{Z_{3}(\log t)^{2/3}(\log\log t)^{1/3}},$
for some positive constants $Z_{1},\,Z_{2},\,Z_{3}$, one usually proves upper bounds for $1/\zeta(s)$ of the order
$\log t,\qquad\frac{\log t}{\log\log t},\qquad(\log t)^{2/3}(\log\log t)^{1/3},$ (1)
respectively. Without making any changes to the zero-free region used, our new method would allow one to refine the results in (1), to
$(\log t)^{11/12},\qquad\frac{(\log t)^{11/12}}{(\log\log t)^{3/4}},\qquad(\log t)^{2/3}(\log\log t)^{1/4},$ (2)
respectively. As far as we are aware, results of these types were not previously known. In this paper, we prove an explicit version of the first bound in (2). The third bound in (2) is the best unconditional bound currently known, and an explicit version of it is given in [19].
2 Main results
Here we address matters of notation before listing some of our main results. We note that the arguments used in this article are largely those of [7].
2.1 Properties of the Riemann zeta function
In what follows, let $s=\sigma+it$, with $\sigma,\,t\in\mathbb{R}$, and define $H:=H_{0}-1/2$, where $H_{0}:=3.\,000\,175\,332\,800\cdot 10^{12}$ is the Riemann height, i.e., the height to which the Riemann hypothesis has been verified (see [22]). Our arguments will consider $t\leq H$ and $t\geq H$ as separate cases.
The classical zero-free region for $\zeta(s)$ states that for $t\geq 2$,
$\zeta(s)\neq 0,\qquad\text{ for }\qquad\sigma\geq 1-\frac{1}{W_{0}\log t}.$
Currently, the best known result is due to [21], where the value $W_{0}=5.558691$ is permissible. We take $W_{0}$ to be this value.
We denote the Riemann hypothesis up to height $T$ by $RH(T)$, by which we mean: for all non-trivial zeros $\rho=\beta+i\gamma$ such that $\zeta(\rho)=0$ and $\gamma\leq T$, we have $\beta=1/2$. Similarly we denote the Riemann hypothesis by $RH$, by which we mean all non-trivial zeros $\rho$ have $\beta=1/2$.
2.2 Bounds on $|\zeta^{\prime}(s)/\zeta(s)|$
Here we present our main results, which are corollaries of the theorems in §4 and §5.
Corollary 1.
Let $W>W_{0}$. In a region
$\sigma\geq 1-\frac{1}{W\log t}$
where $\zeta(\sigma+it)\neq 0$, we have for $t\geq 13$,
$\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq Q\log t,$
where $(W,Q)=(10,71.220)$ is admissible, and other values of $Q$ with corresponding $W$ are presented in Table 3.
Corollary 2.
For $\sigma\geq 1$ and $t\geq 13$, we have
$\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq 24.303\log t.$
Corollary 3.
In the region
$\sigma\geq\sigma_{0}>1/2\qquad\text{ and }\qquad t_{0}\leq t\leq H,$
we have
$\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq Q_{RH(H_{0})}\log t,$
where $(\sigma_{0},\,t_{0},\,Q_{RH(H_{0})})=(0.8,14,23.759)$ is admissible, and other values for $Q_{RH(H_{0})}$ with corresponding values of $\sigma_{0}$ and $t_{0}$ are given in Table 2. Furthermore, by assuming $RH$, the upper bound restriction on $t$ can be removed.
2.3 Bounds on $|1/\zeta(s)|$
Corollary 4.
Let $W>W_{0}$. In a region
$\sigma\geq 1-\frac{1}{W\log t}$
where $\zeta(\sigma+it)\neq 0$, we have for $t\geq 13$,
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq Y\log t,$
where $(W,Y)=(13,1719)$ is admissible, and other values of $Y$ with corresponding $W$ are presented in Table 4.
In addition, we have for $\sigma\geq 1$ and $t\geq 13$,
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq 30.812\log t,$
while for $\sigma=1$ and $t\geq 2$,
$\left|\frac{1}{\zeta(1+it)}\right|\leq 29.388\log t.$
Corollary 5.
Let $W>W_{0}$. In a region
$\sigma\geq 1-\frac{1}{W\log t}$
where $\zeta(\sigma+it)\neq 0$, we have for $t\geq 13$,
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq Y^{\prime}(\log t)^{11/12},$
where $(W,Y)=(13,4904)$ is admissible, and other values of $Y^{\prime}$ with corresponding $W$ are presented in Table 5.
In addition, we have for $\sigma\geq 1$ and $t\geq 13$,
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq 87.725(\log t)^{11/12},$
while for $\sigma=1$ and $t\geq 2$,
$\left|\frac{1}{\zeta(1+it)}\right|\leq 81.900(\log t)^{11/12}.$
3 Some preliminaries
In this section, we present the results used in obtaining our estimates. The argument is largely that of [15] and [7], in which detailed proofs of the core lemmas can be found. At the heart of the method lies the Borel–Carathéodory Theorem (see [15, Theorem 4]), by which the following two lemmas are proven. We remark that Lemma 1 is used in the process of proving Lemma 2.
Theorem 1 (Borel–Carathéodory).
[15, Theorem 14] Let $s_{0}$ be a complex number. Let $R$ be a positive number, possibly depending on $s_{0}$. Suppose that the function $f(s)$ is analytic in a region containing the disc $|s-s_{0}|\leq R$. Let $M$ denote the maximum of $\textup{Re}\,f(s)$ on the boundary $|s-s_{0}|=R$. Then, for any $r\in(0,R)$ and any $s$ such that $|s-s_{0}|\leq r$,
$|f(s)|\leq\frac{2r}{R-r}M+\frac{R+r}{R-r}|f(s_{0})|.$
If in addition $f(s_{0})=0$, then for any $r\in(0,R)$ and any $s$ such that $|s-s_{0}|\leq r$,
$|f^{\prime}(s)|\leq\frac{2R}{(R-r)^{2}}M.$
Lemma 1.
[15, Lemma 15] Let $s_{0}$ be a complex number. Let $r$ and $\alpha$ be positive numbers (possibly depending on $s_{0}$) such that $\alpha<1/2$. Suppose that the function $f(s)$ is analytic in a region containing the disc $|s-s_{0}|\leq r$. Suppose further there is a number $A_{1}$ independent of $s$ such that,
$\left|\frac{f(s)}{f(s_{0})}\right|\leq A_{1},\qquad|s-s_{0}|\leq r.$
Then, for any $s$ in the disc $|s-s_{0}|\leq\alpha r$ we have
$\left|\frac{f^{\prime}(s)}{f(s)}-\sum_{\rho}\frac{1}{s-\rho}\right|\leq\frac{4\log A_{1}}{r(1-2\alpha)^{2}},$
where $\rho$ runs through the zeros of $f(s)$ in the disc $|s-s_{0}|\leq\tfrac{1}{2}r$, counted with multiplicity.
Lemma 2.
[15, Lemma 16] Let $s$ and $s_{0}$ be complex numbers with real parts $\sigma$ and $\sigma_{0}$, respectively. Let $r$, $\alpha$, $\beta$, $A_{1}$ and $A_{2}$ be positive numbers, possibly depending on $s_{0}$, such that $\alpha<1/2$ and $\beta<1$. Suppose that the function $f(s)$ satisfies the conditions of Lemma 1 with $r$, $\alpha$ and $A_{1}$, and that
$\left|\frac{f^{\prime}(s_{0})}{f(s_{0})}\right|\leq A_{2}.$
Suppose, in addition, that $f(s)\neq 0$ for any $s$ in the intersection of the disc $|s-s_{0}|\leq r$ and the right half-plane $\sigma\geq\sigma_{0}-\alpha r$. Then, for any $s$ in the disc $|s-s_{0}|\leq\alpha\beta r$,
$\left|\frac{f^{\prime}(s)}{f(s)}\right|\leq\frac{8\beta\log A_{1}}{r(1-\beta)(1-2\alpha)^{2}}+\frac{1+\beta}{1-\beta}A_{2}.$
The following corollary is useful for proving results in a right-half plane, under $RH(T)$ or $RH$.
Corollary 6.
If in the disc $|s-s_{0}|\leq r$, we have that $f(s)$ is analytic, non-zero, and satisfies $\left|\tfrac{f(s)}{f(s_{0})}\right|\leq A_{1}$, then for $0<\alpha_{0}<1$,
$\left|\frac{f^{\prime}(s)}{f(s)}\right|\leq\frac{2\log A_{1}}{r(1-\alpha_{0})^{2}},\qquad|s-s_{0}|\leq\alpha_{0}r.$
Proof.
We follow a similar proof to that of Lemma 1 (also [15, Lemma 15]). Define $g(s)=\log(f(s)/f(s_{0}))$, where the logarithm branch is determined by $g(s_{0})=0$. Furthermore, since $f(s)\neq 0$ and is analytic in the disc $|s-s_{0}|\leq r$, hence $g(s)$ is analytic there.
By hypothesis, we have $|f(s)/f(s_{0})|\leq A_{1}$ in $|s-s_{0}|\leq r$, so in the same region,
$\textup{Re}(g(s))=\log\left|\frac{f(s)}{f(s_{0})}\right|\leq\log A_{1}.$
To complete the proof, notice that $g^{\prime}(s)=f^{\prime}(s)/f(s)$, and apply Theorem 1 in the two discs $|s-s_{0}|\leq\alpha_{0}r<r$ to obtain
$|g^{\prime}(s)|=\left|\frac{f^{\prime}(s)}{f(s)}\right|\leq\frac{2r\log A_{1}}{(r-\alpha_{0}r)^{2}}=\frac{2\log A_{1}}{r(1-\alpha_{0})^{2}},$
for $|s-s_{0}|\leq\alpha_{0}r$. ∎
Next we require some uniform bounds for $\zeta(s)$. Lemma 4 slightly improves [30, Corollary 1], which is proven using a generalisation of the Phragmén–Lindelöf Principle, due to Trudgian [27, Lemma 3].
Unfortunately, Trudgian’s proof contains an error, the consequence of which is that his result is true only if one includes the restriction $\alpha_{2}=\beta_{2}$ (according to the notation of [27, Lemma 3]). In Lemma 3 below, we present a corrected version which includes this restriction. We refer the reader to [10] for an in-depth discussion on this error and a strategy to recovering the original result [^1] The author thanks Andrew Fiori for bringing this to our attention..
Lemma 3.
Let $a,b,Q$ be real numbers such that $b>a$ and $Q+a>1$. Let $f(s)$ be a holomorphic function on the strip $a\leq\textup{Re}(s)\leq b$ such that
$|f(s)|<C\exp(e^{k|t|})$
for some $C>0$ and $0<k<\tfrac{\pi}{b-a}$. Suppose further that there are $A,B,\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}\geq 0$ such that $\alpha_{1}\geq\beta_{1}$, $\alpha_{2}=\beta_{2}$, and
$|f(s)|\leq\begin{cases}A|Q+s|^{\alpha_{1}}(\log|Q+s|)^{\alpha_{2}}\quad\text{for }\textup{Re}(s)=a;\\ B|Q+s|^{\beta_{1}}(\log|Q+s|)^{\beta_{2}}\quad\text{for }\textup{Re}(s)=b.\end{cases}$
Then for $a\leq\textup{Re}(s)\leq b$, one has
$|f(s)|\leq\Big{(}{A|Q+s|^{\alpha_{1}}|\log(Q+s)|^{\alpha_{2}}}\Big{)}^{\tfrac{b-\textup{Re}(s)}{b-a}}\Big{(}{B|Q+s|^{\beta_{1}}|\log(Q+s)|^{\beta_{2}}}\Big{)}^{\tfrac{\textup{Re}(s)-a}{b-a}}.$
Lemma 4.
Let $h(s)=(s-1)\zeta(s)$, and let $\delta_{r}$ be a positive real number. Furthermore, let $Q_{0}>1/2$ be a number for which
$|h(\tfrac{1}{2}+it)|$ $\leq k_{1}|Q_{0}+\tfrac{1}{2}+it|^{k_{2}+1}(\log|Q_{0}+\tfrac{1}{2}+it|)^{k_{3}}$ (3)
$|h(1+\delta_{r}+it)|$ $\leq k_{4}|Q_{0}+1+\delta_{r}+it|(\log|Q_{0}+1+\delta_{r}+it|)^{k_{3}},$
for all $t$. Then for $\sigma\in[\tfrac{1}{2},1+\delta_{r}]$ and $t\geq t_{0}\geq e$, we have
$|\zeta(s)|\leq$ $\left(k_{4}\right)^{\frac{\sigma-0.5}{0.5+\delta_{r}}}(1+a_{0}(1+\delta_{r},Q_{0},t_{0}))^{k_{3}}\log^{k_{3}}t$ (4)
$\bm{\times}(1+a_{1}(1+\delta_{r},Q_{0},t_{0}))\left(k_{1}(1+a_{1}(1+\delta_{r},Q_{0},t_{0}))^{k_{2}}t^{k_{2}}\right)^{\frac{1+\delta_{r}-\sigma}{0.5+\delta_{r}}},$
where
$a_{0}(\sigma,Q_{0},t)=\frac{\sigma+Q_{0}}{2t^{2}\log t}+\frac{\pi}{2\log t}+\frac{\pi(\sigma+Q_{0})^{2}}{4t\log^{2}t},\qquad a_{1}(\sigma,Q_{0},t)=\frac{\sigma+Q_{0}}{t}.$
In addition, the right-hand side of (4) is decreasing in $\sigma$, provided that $k_{i}\geq 0$, and the following hold:
$t\geq\left(\frac{k_{4}}{k_{1}(1+a_{1}(1+\delta_{r},Q_{0},t_{0}))^{k_{2}}}\right)^{1/k_{2}}.$ (5)
Proof.
The proof is analogous to that of [30, Lemma 2.1], through a direct application of Lemma 3 with $a=1/2$ and $b=1+\delta_{r}$, except that we are more precise in estimating (4) by retaining the dependence on $\sigma$ in the exponents of $1+a_{1}(\sigma,Q_{0},t)$ and $1+a_{0}(\sigma,Q_{0},t)$.
Note that different to the proof of [30, Lemma 2.1], we use only two lines, $\sigma=1/2$ and $\sigma=1+\delta_{r}$, omitting the need for a bound at $\sigma=1$. This is because we intend to find the maximum bound over $\sigma\in[1/2,1+\delta_{r}]$, which we show occurs at $\sigma=1/2$, as our bound is decreasing in $\sigma$. The decreasing condition can be verified by checking when the derivative of the right-hand side of (4) is negative, and this is what gives rise to the condition in (5). Additionally, we found that further splitting the interval complicates (5) and raises the threshold of $t$ for the lemma to be valid. ∎
Corollary 7.
For $\sigma\in[\tfrac{1}{2},1+\delta_{r}]$ and $t\geq t_{0}\geq e$, we have
$|\zeta(s)|\leq 0.618(1+a_{0}(1+\delta_{r},1.31,t_{0}))(1+a_{1}(1+\delta_{r},1.31,t_{0}))^{7/6}t^{1/6}\log t,$
provided that
$t^{1/6}\log t\geq\frac{\zeta(1+\delta_{r})}{0.618(1+a_{1}(1+\delta_{r},1.31,t_{0}))^{1/6}},$ (6)
where $\delta_{r}$, $a_{0}(\sigma,1.31,t)$, and $a_{1}(\sigma,1.31,t)$ are defined as in Lemma 4.
Proof.
Recall that Hiary, Patel, and Yang [16] have $|\zeta(1/2+it)|\leq 0.618t^{1/6}\log t$, for $t\geq 3$, while a quick calculation on Mathematica shows that $|\zeta(1/2+it)|\leq 1.461$ for $|t|<3$ (see also [14]). Meanwhile, one has by $|\zeta(s)|\leq\zeta(\sigma)$ that for all $t$,
$|\delta_{r}+it||\zeta(1+\delta_{r}+it)|\leq\frac{\zeta(1+\delta_{r})}{\log t}|\delta_{r}+it|\log t.$
Therefore, it can be verified that both bounds in (3) are met.
We now refer to Lemma 4, taking the maximum of (4) at $\sigma=1/2$. It has been shown in [31, p. 15] that $Q_{0}=1.31$ is permissible in (3) together with
$(k_{1},k_{2},k_{3},k_{4},Q_{0})=\left(0.618,\tfrac{1}{6},1,\tfrac{\zeta(1+\delta_{r})}{\log t},1.31\right),$
hence we are done. ∎
Lemma 5.
For every $|t|\geq 3$, we have
$|\zeta(1+it)|\leq 58.096(\log|t|)^{2/3}.$
Proof.
We use same argument in the first paragraph of [28, p. 260], which modifies Ford’s proof in [11, Lemma 7.3]. The only difference is that we now apply the argument with parameters $C$ and $D$ found in the proof of Theorem $1.1$ of [2, §5], which itself is an improvement on Ford’s result. This proves the lemma. ∎
Lemma 6.
For $t\geq t_{0}\geq 3$ and $\sigma\geq 1$, we have
$|\zeta(\sigma+it)|\leq C_{3}(\log t)^{2/3},$
with $a_{0}$ defined as in Lemma 4, and
$C_{3}=58.096\sqrt{1+9t_{0}^{-2}}\left(1+a_{0}(2,1,t_{0})\right)^{2/3}.$ (7)
Proof.
The proof is similar to that of [15, Lemma 7]. First, to show that the result holds for $1\leq\sigma\leq 2$, we use the Phragmén–Lindelöf Principle, i.e., Lemma 3, on the holomorphic function
$f(s)=(s-1)\zeta(s).$
To this end, on the $1$-line we have
$|f(1+it)|=|t|\,|\zeta(1+it)|\leq 58.096\,|2+it|(\log|2+it|)^{2/3}.$
This inequality is verified numerically via Mathematica for $|t|<3$, and is a consequence of Lemma 5 for $|t|\geq 3$.
Also, on the line $s=2+it$, the trivial bound $|\zeta(s)|\leq\zeta(\sigma)$ gives
$|f(2+it)|\leq|1+it|\,\zeta(2)<58.096|3+it|(\log|3+it|)^{2/3},$
for all real $t$. Now apply Lemma 3 with the following parameter values: $a=1$, $b=2$, $Q=1$, $\alpha_{1}=\beta_{1}=1$, $\alpha_{2}=\beta_{2}=2/3$, and $A=B=58.096$. We obtain
$|f(s)|\leq 58.096|1+s||\log(1+s)|^{2/3},$ (8)
for $1\leq\sigma\leq 2$. Moreover, for the same range of $\sigma$, we have the following estimates:
$\begin{split}\frac{|1+s|}{|s-1|}&\leq\frac{\sqrt{9+t^{2}}}{t}\leq\sqrt{1+9t_{0}^{-2}},\\ |\log(1+s)|&\leq\left(1+a_{0}(2,1,t_{0})\right)\log t,\end{split}$
where $a_{0}$ is defined as in Lemma 4. The second inequality above is derived using the same argument as that in [30, §2].
Thus for $1\leq\sigma\leq 2$ and $t\geq t_{0}$, the inequality (8) implies that
$|\zeta(\sigma+it)|$ $\leq 58.096\sqrt{1+9t_{0}^{-2}}\left(1+a_{0}(2,1,t_{0})\right)^{2/3}\log^{2/3}t\leq C_{3}\log^{2/3}t,$
as desired, where $C_{3}$ is defined in (7).
Finally, we address the case where $\sigma\geq 2$. Here since $\sigma>1$, we apply the trivial bound
$|\zeta(s)|\leq\zeta(\sigma)\leq\frac{\sigma}{\sigma-1},$ (9)
where $\sigma/(\sigma-1)$ is decreasing in $\sigma$. Then for $\sigma\geq 2$, it is easy to verify that
$|\zeta(\sigma+it)|\leq 2\leq C_{3}(\log t)^{2/3}$
for all $t\geq t_{0}$. ∎
Lemma 7.
[24, Lemma 4] In the region $1/2<\sigma\leq 3/2$ and $0\leq t\leq 2\exp(e^{2})$, we have
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq\frac{4}{\sigma-(1/2)}.$
In the remainder of this section, we present some estimates concerning the Riemann zeta function, including some for $\sigma>1$, which are better than the trivial ones in some finite range.
We would like to point out that Lemmas 10 and 11, as well as Corollary 10, do not contribute to the main results of this paper, but are included as they are potentially useful estimates for anyone seeking to do explicit computations involving $\zeta(s)$.
Similarly, Lemma 8 and Corollary 9, which can both be viewed as an extension of Corollary 8 inside the critical strip, are not actually required for the main purposes of this paper. However, we include them here as the proofs are similar to that of Corollary 8, and at the same time we provide a brief discussion on their usefulness. Lemma 8 is an improvement on [17, Proposition 5.3.2], which proves a simlar result.
Lemma 8.
Let $W>0$ be a real constant. Then in the region
$1-\frac{1}{W\log t}\leq\sigma\leq\sigma_{1},$
we have for $t\geq t_{0}\geq 3$ and any $\eta\in[2/t_{0},1-1/t_{0}]$ that
$|\zeta(\sigma+it)|\leq e^{1/W}\log t+C_{0}(W,\sigma_{1},t_{0},\eta)$
where
$C_{0}(W,\sigma_{1},t_{0},\eta)=$ $e^{1/W}\Bigg{(}\log\left(\eta+\frac{1}{t_{0}}\right)+\gamma+\frac{1}{12\eta^{2}}$
$+\frac{1}{t_{0}}\left(\frac{1}{6\eta^{2}}+\frac{1}{\eta}+1\right)+\frac{16(\sigma_{1}^{2}+2\sigma_{1}-1)+3}{192\eta^{2}t_{0}^{2}}\Bigg{)}.$
Proof.
We will use the following representation for $\zeta(s)$ for $N\geq 2$,
$\zeta(s)=\sum_{n=1}^{N-1}\frac{1}{n^{s}}+\frac{1}{2N^{s}}+\frac{N^{1-s}}{s-1}+\frac{s}{12N^{s+1}}-\frac{s(s+1)}{2}\int_{N}^{\infty}\frac{\varphi^{*}(u)}{u^{s+2}}\text{d}u,$ (10)
where $\varphi^{*}(u)$ is the periodic function obtained by extending the polynomial $\varphi(u)=u^{2}-u+1/6$ on the interval $[0,1]$ (for full details see e.g., [1, Equation (8)]). Note that (10) holds for $\sigma>0$ by analytic continuation (see [26, Theorem 3.5]).
In the region of $\sigma$ considered, we have
$|N^{-s}|=N^{-\sigma}\leq\exp\left(-\Big{(}1-\frac{1}{W\log t}\Big{)}\log N\right)=N^{-1}\exp\left(\frac{\log N}{W\log t}\right).$ (11)
This bound can be used throughout (10). The first term in (10) can also be bounded using the estimate in [12, Lemma 2.1], so for $N\geq 2$,
$\left|\sum_{n=1}^{N-1}\frac{1}{n^{s}}+\frac{1}{2N^{s}}\right|$ $=\left|\sum_{n=1}^{N}\frac{1}{n^{s}}-\frac{1}{2N^{s}}\right|$
$\leq\exp\left(\frac{\log N}{W\log t}\right)\left(\sum_{n=1}^{N}\frac{1}{n}+\frac{1}{2N}\right)$
$\leq\exp\left(\frac{\log N}{W\log t}\right)\left(\log N+\gamma+\frac{1}{N}-\frac{1}{12N^{2}}+\frac{1}{64N^{4}}\right).$
The term $N^{1-s}/(s-1)$ will be bounded by $t^{-1}\exp\left(\frac{\log N}{W\log t}\right)$. The integral in (10) can be evaluated using the bound $|\varphi^{*}(u)|\leq 1/6$, so
$\int_{N}^{\infty}\frac{|\varphi^{*}(u)|}{u^{\sigma+2}}\text{d}u\leq\frac{1}{6(\sigma+1)}\frac{1}{N^{\sigma+1}}\leq\frac{1}{6}\exp\left(\frac{\log N}{W\log t}\right)\frac{1}{N^{2}}.$
This implies
$\left|\frac{s}{12N^{s+1}}-\frac{s^{2}+s}{2}\int_{N}^{\infty}\frac{\varphi^{*}(u)}{u^{s+2}}\text{d}u\right|$ $\leq\exp\left(\frac{\log N}{W\log t}\right)\left(\frac{2|s|}{12N^{2}}+\frac{|s^{2}|}{12N^{2}}\right)$
$\leq\exp\left(\frac{\log N}{W\log t}\right)\left(\frac{t+\sigma_{1}}{6N^{2}}+\frac{t^{2}+\sigma_{1}^{2}}{12N^{2}}\right).$
We thus have that $|\zeta(s)|$ is upper bounded by
$\exp\left(\frac{\log N}{W\log t}\right)\left(\log N+\gamma+\frac{1}{N}+\frac{1}{64N^{4}}+\frac{1}{t}+\frac{t^{2}+\sigma_{1}^{2}+2(t+\sigma_{1})-1}{12N^{2}}\right).$
It remains to choose $N$. We can minimise the expression by taking $N=\lfloor\eta t\rfloor+1$ for some constant $\eta\in[2/t_{0},1-1/t_{0}]$. This will make
$\exp\left(\frac{\log N}{W\log t}\right)\leq\exp\left(\frac{1}{W}\right)$
for all $t\geq t_{0}$. The bound then becomes
$|\zeta(s)|$ $\leq e^{1/W}\left(\log t+\log(\eta+t^{-1})+\gamma+\frac{1}{t}\left(\frac{1}{\eta}+1\right)+\frac{1}{64\eta^{2}t^{2}}\right$.
$\hskip 142.26378pt+\left.\frac{1}{12\eta^{2}}+\frac{1}{6\eta^{2}t}+\frac{\sigma_{1}^{2}+2\sigma_{1}-1}{12\eta^{2}t^{2}}\right)$
$\leq e^{1/W}\log t+C_{0}(W,\sigma_{1},t_{0},\eta)$
for $t\geq t_{0}$. ∎
We also provide a corollary, specialised for $\sigma\geq 1$, which sharpens the above estimates.
Corollary 8.
Let $s=\sigma+ikt$ with $1\leq\sigma\leq\sigma_{1}\leq 2$, $t\geq t_{0}\geq 3$, and a fixed $k\geq 1$. For $0<\eta/k<t_{0}$, we have
$|\zeta(\sigma+ikt)|<\log(kt)+C(\sigma_{1},t_{0},k,\eta),$
where
$C(\sigma_{1},t_{0},k,\eta)=$ $\log\left(\frac{1}{\eta}+\frac{1}{kt_{0}}\right)+\gamma+\frac{\eta^{4}}{64(kt_{0})^{4}}$ (12)
$+\frac{12+(c_{1}(\sigma_{1},t_{0},k)-(kt_{0})^{-1})\eta^{2}}{12kt_{0}}+\frac{c_{2}(\sigma_{1},t_{0},k)\eta^{2}}{24},$
with $\gamma$ denoting Euler’s constant, and
$c_{1}(\sigma_{1},t_{0},k)$ $=\left(\frac{\sigma_{1}^{2}}{(kt_{0})^{2}}+1\right)^{1/2},$
$c_{2}(\sigma_{1},t_{0},k)$ $=c_{1}(\sigma_{1},t_{0},k)\left(\frac{(\sigma_{1}+1)^{2}}{(kt_{0})^{2}}+1\right)^{1/2}.$
Proof.
The proof follows the argument of [3, Proposition A.1], originally due to Backlund [1], and extends their result to $t<500$. Note that here we have instead the choice of $N=\lfloor kt/\eta\rfloor+1$. A slightly better estimate of the harmonic sum from [12, Lemma 2.1] was also used. We omit the proof since the process is almost identical to that of Lemma 8. ∎
To use Lemma 8, we want to choose a $t_{0}$ and optimise over $\eta$. For the smallest $t_{0}=3$, we can only use $\eta=2/3$. However, as $t_{0}\rightarrow\infty$, the optimal $\eta$ quickly approaches $1/\sqrt{6}\approx 0.4082$. We also have the following corollary, which is not required for our purposes, but is useful to have for other applications.
Corollary 9.
In the region
$1-\frac{1}{W_{0}\log t}\leq\sigma\leq 1,$ (13)
we have for $t\geq t_{0}\geq 3$,
$|\zeta(s)|\leq 1.1971\log t+B_{1},$ (14)
where Table 1 gives values of $B_{1}$ for select $t_{0}$.
Proof.
For a fixed $t_{0}$ and $\sigma_{1}=1$, in Lemma 8 optimise over $\eta$ to find the minimal $B_{1}$. ∎
   $t_{0}$    $B_{1}$    $\eta$
   33    $2.1173$    $2/3$
   $10^{3}$    $0.2254$    $0.41$
   $10^{12}$    $0.2171$    $0.41$
Table 1: Values of $B_{1}$ for Corollary 9 with the corresponding optimised $\eta$ from Lemma 8, where $\sigma_{1}=1$.
Corollary 9 can be used in conjunction with other estimates for $\zeta(s)$ in the critical strip, such as Ford’s Theorem 1 in [11] or Trudgian’s Lemma 2.1 [30]. Ford’s bound, while asymptotically better, is only sharper than bounds of the form $\zeta(s)\ll\log t$ for very large $t$. For $\sigma\in[1-1/(W_{0}\log t),1]$ and $3\leq t\leq 10^{10^{5}}$, Corollary 9 is an improvement to [11, Theorem 1]. This makes it useful for results over smaller $t$.
It is not uncommon to degrade asymptotically sharper results for $\zeta(s)$ to $|\zeta(s)|\leq C\log t$ for the sake of a smaller main-term constant $C$. Corollary 9 now directly offers a better constant. For instance, over $\sigma\in[1-1/(W_{0}\log t),1]$ and $t\geq 3$ our result implies
$|\zeta(s)|\leq 3.13\log t,$
whereas degrading a Richert-type bound like Ford’s [11, Theorem 1] would give
$|\zeta(s)|\leq 102\log t,$
and a bound obtained via the Phragmén–Lindelöf principle like (2.4) in [30, Lem. 2.1] using estimates on the $1/2$-line (see Hiary, Patel, and Yang [16]) and estimates on the $1$-line (see Backlund [1]) would yield
$|\zeta(s)|\leq 364\log t.$
We also remark that although Lemma 6 is asymptotically superior to Corollary 8, one would expect that the former beats the latter only for a sufficiently large $t$; approximately $t\geq\exp(58^{3})\geq 10^{84736}$, say. This can be improved by lowering the constant in Lemma 5, and correspondingly $C_{3}$.
Next we have some bounds on $\zeta(\sigma)$ for real $\sigma>1$. These are useful in their own right, but we show they can be used to obtain good bounds on $\zeta^{\prime}(\sigma)/\zeta(\sigma)$ in a similar region.
Lemma 9.
[23, Lemma 5.4] Let $\gamma$ denote Euler’s constant. For $\sigma>1$ we have
$|\zeta(\sigma+it)|\leq\zeta(\sigma)\leq\frac{e^{\gamma(\sigma-1)}}{\sigma-1}.$
Lemma 10.
Let $\gamma_{n}$ be the Laurent–Stieltjes constants (with $\gamma:=\gamma_{0}$). For integers $k$ such that $1\leq k\leq 1500$, and real $\sigma$ satisfying $1<\sigma\leq\sigma_{k}<3$, we have
$\zeta(\sigma)\geq\frac{1}{\sigma-1}+\gamma+\phi_{0}(\sigma,k),$
where
$\phi_{0}(\sigma,k)=\sum_{n=1}^{k}\frac{(-1)^{n}(\sigma-1)^{n}}{n!}\gamma_{n}+\frac{1}{(\sigma-3)}\left(\frac{\sigma-1}{2}\right)^{k+1}\geq 0.$ (15)
In addition, we note some admissible pairs:
$(\sigma_{k},k)=(1.45,1),\,(2,3),\,(2.51,10),\,(2.97,500),\,(2.98,1000),\,(2.99,1500).$
Proof.
To prove the first assertion of the lemma, we have from the Laurent series expansion (about $\sigma=1$) of the zeta function
$\zeta(\sigma)$ $=\frac{1}{\sigma-1}+\gamma+\sum_{n=1}^{k}\frac{(-1)^{n}(\sigma-1)^{n}\gamma_{n}}{n!}+\sum_{n=k+1}^{\infty}\frac{(-1)^{n}(\sigma-1)^{n}\gamma_{n}}{n!}$ (16)
$\geq\frac{1}{\sigma-1}+\gamma+\sum_{n=1}^{k}\frac{(-1)^{n}(\sigma-1)^{n}\gamma_{n}}{n!}-\left|\sum_{n=k+1}^{\infty}\frac{(-1)^{n}(\sigma-1)^{n}\gamma_{n}}{n!}\right|.$
We have the bounds on $\gamma_{n}$ due to Lavrik [18]:
$|\gamma_{n}|\leq\frac{n!}{2^{n+1}},\quad\quad n=1,2,3,\ldots,$ (17)
hence
$\left|\sum_{n=k+1}^{\infty}\frac{(-1)^{n}(\sigma-1)^{n}\gamma_{n}}{n!}\right|\leq\sum_{n=k+1}^{\infty}\frac{(\sigma-1)^{n}}{2^{n+1}}=-\frac{1}{(\sigma-3)}\left(\frac{\sigma-1}{2}\right)^{k+1}$
by the geometric series with $|1-\sigma|<2$. Taking the lower bound from this, one can then verify that $\phi_{0}(\sigma,k)\geq 0$ for $1<s\leq\sigma_{k}$ by calculating the first $k$ terms. ∎
It is likely that $\phi_{0}(\sigma,k)$ is always positive for any $\sigma>1$, although proving that would require a different approach, perhaps one akin to [23, Lemma 5.3]. Our method in Lemmas 10 and 11, while simpler, is restricted by bounds on $|\gamma_{n}|$ which loses some information about sign changes. It is worth noting that bounds sharper than (17) are available, but are more complicated to work with.
Lemma 11.
Let the hypotheses of Lemma 10 be fulfilled. Then for $\sigma$ real, we have
$\zeta^{\prime}(\sigma)\leq\frac{1}{(\sigma-1)^{2}}+\frac{1}{(\sigma-3)^{2}},\qquad 1<\sigma<3,$
and
$\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq\frac{-\zeta^{\prime}(\sigma)}{\zeta(\sigma)}\leq\frac{\phi_{1}(\sigma,k)}{\sigma-1},\qquad 1<\sigma\leq\sigma_{k},$
where
$\phi_{1}(\sigma,k)=\frac{(\sigma-3)^{2}+(\sigma-1)^{2}}{(\sigma-3)^{2}(1+(\sigma-1)(\phi_{0}(\sigma,k)+\gamma))}.$ (18)
Proof.
The first inequality of the second assertion is trivial. The subsequent inequality is proved by differentiating (16) and repeating a process similar to the proof of Lemma 10 (not splitting the sum for brevity), to obtain
$-\zeta^{\prime}(\sigma)$ $\leq\frac{1}{(\sigma-1)^{2}}+\sum_{n=1}^{\infty}\frac{n(\sigma-1)^{n-1}}{2^{n+1}}$
$=\frac{1}{(\sigma-1)^{2}}+\frac{1}{2(\sigma-1)}\sum_{n=1}^{\infty}n\left(\frac{\sigma-1}{2}\right)^{n}=\frac{1}{(\sigma-1)^{2}}+\frac{1}{(\sigma-3)^{2}}$
for $|1-\sigma|<2$ by identification with the polylogarithm $\textup{Li}_{-1}(z)=z/(1-z)^{2}$. Dividing by the bound in Lemma 10 proves the result. ∎
These bounds are good for $\sigma$ close to $1$. For instance, we have that
$\phi_{1}(\sigma,3)<1\qquad\text{for}\qquad\sigma\leq 1.83,$ (19)
which is an improvement on Delange’s result [8] for a fixed $\sigma\leq 1.65$. Although the bounds get sharper as $k$ increases, in practice, taking $k\leq 10$ is more than sufficient for small $\sigma$ and any savings gained by subsequent increases in $k$ would be after the fifth decimal place.
Corollary 10.
Let $\phi_{1}(\sigma,k)$ be defined as in (18). For $1<\sigma_{0}\leq\sigma\leq 1.5$, we have
$\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq\frac{-\zeta^{\prime}(\sigma)}{\zeta(\sigma)}\leq\frac{\phi_{2}(\sigma_{0})}{\sigma-1},$
where $\phi_{2}(\sigma_{0})=\max\left\{\phi_{1}(\sigma_{0},10),\,0.852\right\}$.
Proof.
For $k=10$, we have from Lemma 16 that $\phi_{0}(\sigma,10)\geq 0$ in our stated range of $\sigma$. It can also be verified that $\phi_{1}(\sigma,10)$ is decreasing in the region $1<\sigma\leq 1.48$, and therefore $\phi_{1}(\sigma,10)\leq\phi_{1}(\sigma_{0},10)$.
To complete the proof, verify that $\phi_{1}(\sigma,10)\leq 0.852$ in the region $1.48\leq\sigma\leq 1.5$. ∎
Finally, we have a lower bound on $|\zeta(s)|$, for $\sigma>1$, which is better than the trivial bound $|\zeta(s)|\geq\zeta(2\sigma)/\zeta(\sigma)$ over certain ranges of $\sigma$ (for more information, see [20]). This is the $\log t$ analogue of [7, Lemma 8].
Lemma 12.
Let $\kappa,\,\sigma_{1}$ be positive parameters, and $\gamma$ be Euler’s constant. Then for $t\geq t_{0}\geq 3$ and $1+\kappa\leq\sigma_{1}\leq 2$, we have
$\left|\frac{1}{\zeta(1+\kappa+it)}\right|\leq V(\kappa,\sigma_{1},t_{0},\eta)\log t$
with
$V(\kappa,\sigma_{1},t_{0},\eta):=\left(\frac{\exp\left(\gamma\kappa\right)}{\kappa\log t_{0}}\right)^{3/4}\left(1+\frac{\log 2+C(\sigma_{1},t_{0},2,\eta)}{\log t_{0}}\right)^{1/4},$
where $\eta$ and $C(\sigma_{1},t_{0},k,\eta)$ are defined in Corollary 8.
Proof.
Since $\sigma>1$, we use the classical non-negativity argument involving the trigonometric polynomial $2(1+\cos\theta)^{2}=3+4\cos\theta+\cos 2\theta$ (see [26, § 3.3] and [20]). That is, we use $\zeta^{3}(\sigma)|\zeta^{4}(\sigma+it)\zeta(\sigma+2it)|\geq 1$, which implies
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq|\zeta(\sigma)|^{3/4}|\zeta(\sigma+2it)|^{1/4}.$ (20)
Taking $\sigma=1+\kappa$ in (20) gives
$\left|\frac{1}{\zeta(1+\kappa+it)}\right|\leq|\zeta(1+\kappa)|^{3/4}|\zeta(1+\kappa+2it)|^{1/4}.$ (21)
The two factors on the right-hand side can be bounded with Lemma 9 and Corollary 8, respectively, with the aid of the bound
$\log 2t$ $\leq\left(1+\frac{\log 2}{\log t_{0}}\right)\log t.$
This proves the lemma. ∎
4 Estimates for $|\zeta^{\prime}(s)/\zeta(s)|$
4.1 Assuming a partial Riemann hypothesis
In this section, we assume $RH(T)$, and estimate $|\zeta^{\prime}(s)/\zeta(s)|$ in a right-half plane: $\sigma>1/2+\epsilon$, $t_{0}\leq t\leq T$.
Theorem 2.
Assume $RH(T)$. Let $\alpha_{0}$ and $\epsilon$ be parameters, such that $0<\epsilon\leq 1/2$ and $0<\alpha_{0}<1$. In the region
$\sigma\geq 1+\epsilon-\frac{\alpha_{0}}{2}\quad\text{and}\quad 3\leq t_{0}\leq t\leq T-\frac{1}{2},$
we have
$\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq Q_{RH(T)}\log t,$
with
$Q_{RH(T)}$ $=\max\Bigg{\{}\frac{4}{(1-\alpha_{0})^{2}}\Bigg{(}\frac{1}{6}+2B_{t_{0}}+\frac{\log A_{3}}{\log t_{0}}\Bigg{)}\,,\,\frac{1}{(\epsilon+(\alpha_{0}/2))\log t_{0}}\Bigg{\}},$ (22)
where $A_{3}$ is defined in (4.1), and
$B_{t_{0}}=\begin{cases}\frac{1}{e}&\text{ if }t_{0}\leq e^{e},\\ \frac{\log\log t_{0}}{\log t_{0}}&\text{ if }t_{0}>e^{e},\end{cases}$ (23)
provided that
$t^{1/6}\log t\geq\frac{\zeta(\frac{3}{2}+\epsilon)}{0.618(1+a_{1}(\frac{3}{2}+\epsilon,1.31,t_{0}))^{1/6}}$ (24)
is satisfied, with $a_{1}(\sigma,1.31,t)$ defined as in Lemma 4.
Furthermore, by assuming $RH$ instead of $RH(T)$, the upper bound restriction on $t$ can be removed.
Proof.
We will first prove the result for $T_{0}$ and then make the substitution to $t$ at the end. The idea is to construct discs $|s-s_{0}|\leq 1/2$ which have origins slightly to the right of the $1$-line, centred at $s_{0}=\sigma_{0}+iT_{0}$, where
$\sigma_{0}=1+\epsilon,\qquad 3\leq t_{0}\leq T_{0}\leq T-1/2,$
and $\epsilon\leq 1/2$ is a parameter to be determined later. We then extend their radii into the critical strip, so that Corollary 6 applied with
$f(s):=\zeta(s),\qquad s_{0}=\sigma_{0}+iT_{0},\qquad r=1/2,$
gives us the desired estimates. Due to $RH(T)$, $\zeta(s)$ is analytic and non-zero in the discs.
In order to do this, we first need to determine $A_{1}$. Let $V(\epsilon,\sigma_{1},t_{0},\eta)$ be defined as in Lemma 12, with $\sigma_{1}$ and $\eta$ as parameters. Then by Corollary 7 with $\delta_{r}=r+\epsilon$, and Lemma 12 with $\kappa=\epsilon$, we have for $1/2\leq\sigma\leq\sigma_{0}+r$ and each $T_{0}\in[T_{0}-r,T_{0}+r]$,
$\left|\frac{\zeta(\sigma+iT_{0})}{\zeta(\sigma_{0}+iT_{0})}\right|\leq A_{1}:=A_{3}T_{0}^{1/6}(\log T_{0})^{2}$ (25)
where
$A_{3}:=0.618$ $(1+a_{0}(\sigma_{0}+r,1.31,t_{0}))(1+a_{1}(\sigma_{0}+r,1.31,t_{0}))^{7/6}$
$\bm{\cdot}\left(1+\frac{r}{t_{0}}\right)^{1/6}\left(1+\frac{\log\left(1+\frac{r}{t_{0}}\right)}{\log t_{0}}\right)V(\epsilon,\sigma_{1},t_{0},\eta);$ (26)
provided (6) is satisfied.
Now choosing a fixed positive $\alpha_{0}<1$, we have by Corollary 6,
$\left|\frac{\zeta^{\prime}(\sigma+iT_{0})}{\zeta(\sigma+iT_{0})}\right|$ $\leq\frac{4}{(1-\alpha_{0})^{2}}\left(\frac{1}{6}\log T_{0}+2\log\log T_{0}+\log A_{3}\right)$
$\leq\frac{4}{(1-\alpha_{0})^{2}}\left(\frac{1}{6}+2B_{t_{0}}+\frac{\log A_{3}}{\log t_{0}}\right)\log T_{0},$ (27)
which holds for
$1+\epsilon-\frac{\alpha_{0}}{2}\leq\sigma\leq 1+\epsilon+\frac{\alpha_{0}}{2}.$
Note that $B_{t_{0}}$ is defined in (23), and is utilised since $\log\log t/\log t$ attains its maximum at $t=e^{e}$, after which it is decreasing.
To obtain a bound that is valid also for $\sigma\geq 1+\epsilon+\alpha_{0}/2$, we use the trivial bound for all $\sigma>1$ [13, Lemma 70.1]:
$\left|\frac{\zeta^{\prime}(s)}{\zeta(s)}\right|\leq-\frac{\zeta^{\prime}(\sigma)}{\zeta(\sigma)}<\frac{1}{\sigma-1}$ (28)
with $\sigma=1+\epsilon+\alpha_{0}/2$, and then take the maximum of (4.1) and (28).
Finally, making the substitution from $T_{0}$ to $t$ proves the result. ∎
Assuming $RH(H_{0})$ allows us to prove Corollary 3, which holds for $t\leq H$.
Proof of Corollary 3.
For $t\geq t_{0}$ and $\sigma\geq\sigma_{0}$, in Theorem 2 set $\sigma_{0}=1+\epsilon-\frac{\alpha_{0}}{2}$. Then fixing $t_{0}$ and $\sigma_{0}$, optimise (22) over $\epsilon$ (which also determines $\alpha_{0}$), and also over $\sigma_{1},\,\eta$ (due to the quantity $V$ in $A_{3}$), while ensuring that the restrictions set in Theorem 2 (and by extension Lemma 12 and Corollary 8), hold. The value $t_{0}$ is chosen by finding the smallest value of $t$ that satisfies (24). ∎
$\sigma_{0}$ $t_{0}$ $Q_{RH(H_{0})}$ $\epsilon$ $\sigma_{1}$ $\eta$
$0.51$ $15$ $31447$ $0.000462$ $1.965754$ $2.608856$
$0.55$ $14$ $1086$ $0.002766$ $1.901324$ $3.329589$
$0.6$ $14$ $249.260$ $0.006041$ $1.795116$ $3.162609$
$0.65$ $14$ $104.992$ $0.009573$ $1.846979$ $3.195987$
$0.7$ $14$ $56.739$ $0.013287$ $1.797626$ $3.198866$
$0.75$ $14$ $35.162$ $0.017144$ $1.155376$ $2.868261$
$0.8$ $14$ $23.759$ $0.021126$ $1.392644$ $3.173843$
$0.85$ $14$ $17.050$ $0.025202$ $1.763188$ $3.402204$
$0.9$ $13$ $13.123$ $0.029430$ $1.227811$ $3.283187$
$0.95$ $13$ $10.175$ $0.033682$ $1.227876$ $3.283802$
11 $13$ $8.101$ $0.037999$ $1.889284$ $3.054339$
Table 2: Values for $Q_{RH(H_{0})}$ and $\sigma\geq\sigma_{0}$ in Corollary 3, with corresponding $\epsilon,\,\sigma_{1},\,\eta$ (all rounded to six decimal places). Each entry is valid for $t_{0}\leq t\leq H$.
We note that [5, Theorem 1.1] implies that
$\left|\frac{\zeta^{\prime}(1+it)}{\zeta(1+it)}\right|\leq 0.639\log t$
for $10^{6}\leq t\leq 2.99997\cdot 10^{12}$. This result allows us to show that
$\left|\frac{\zeta^{\prime}(1+it)}{\zeta(1+it)}\right|\leq 40.440\log\log t$ (29)
holds unconditionally for $e^{e}\leq t\leq 2.99997\cdot 10^{12}$.
To prove this, recall from §2.1 that $RH$ has been verified up to $3.000175332800\cdot 10^{12}$. Now note that $\log t/\log\log t$ is increasing for $t\geq e^{e}$. Then repeating the process in the proof of Corollary 3, for $e^{e}\leq t\leq 10^{6}$, we obtain
$\left|\frac{\zeta^{\prime}(1+it)}{\zeta(1+it)}\right|\leq 7.686\log t\leq 40.440\log\log t,$
with rounded parameters $(\epsilon,\sigma_{1},\eta)=(0.037816,1.395842,3.177141)$. The second inequality above uses the fact that $t$ attains its maximum at $10^{6}$.
We remark that (29) is an unconditional result due to the partial verification of $RH$. While this is nowhere near as sharp as the bounds in [5], it nevertheless holds for a wider range and might be useful for smaller values of $t$.
If desired, one can extend the range for which the corollary is unconditional beyond this (albeit still in a finite range $t_{0}\leq t\leq T$), at the expense of a worse leading constant. This will prove a poor man’s version of [5, Theorem 1.1], but with a simpler method. As mentioned in §1, the range of validity is wider but the constants are worse. This extension requires a modification in the proof of Theorem 2, by fixing $r=\epsilon+1/W\log T$, utilising a classical zero-free region, and also using Lemma 8 in place of Corollary 7 when estimating $A_{1}$ in (25).
4.2 Above the Riemann height
$W$ $Q$ $\beta$ $\epsilon_{1}$ $\sigma_{1}$ $\eta$
$5.559$ $928465$ $0.999980$ $0.012554$ $1.756408$ $1.050104$
$5.56$ $170199$ $0.999890$ $0.012551$ $1.937703$ $2.228745$
$5.5601$ $157356$ $0.999881$ $0.012563$ $1.393177$ $4.286069$
$5.5602$ $146346$ $0.999872$ $0.012587$ $1.710791$ $1.367807$
$5.5603$ $136750$ $0.999863$ $0.012582$ $1.776753$ $1.141731$
$5.5604$ $128379$ $0.999854$ $0.012712$ $1.889117$ $2.700506$
$5.5605$ $120884$ $0.999845$ $0.012614$ $1.711461$ $3.302948$
$5.561$ $93736$ $0.999800$ $0.012690$ $1.339371$ $2.627981$
$5.562$ $64686$ $0.999710$ $0.012842$ $1.299159$ $2.599102$
$5.563$ $49364$ $0.999620$ $0.012699$ $1.427545$ $1.952944$
$5.564$ $39919$ $0.999531$ $0.012747$ $1.557141$ $2.691489$
$5.565$ $33530$ $0.999441$ $0.013347$ $1.715049$ $4.337899$
$5.566$ $28887$ $0.999351$ $0.013295$ $1.234376$ $4.393614$
$5.567$ $25383$ $0.999262$ $0.013634$ $1.973530$ $3.744506$
$5.568$ $22625$ $0.999172$ $0.013225$ $1.814250$ $2.845626$
$5.569$ $20414$ $0.999082$ $0.013180$ $1.641377$ $1.985693$
$5.57$ $18592$ $0.998993$ $0.012868$ $1.322806$ $4.407476$
$5.575$ $12870$ $0.998545$ $0.012963$ $1.920818$ $3.555767$
$5.58$ $9863$ $0.998102$ $0.019214$ $1.690970$ $2.264745$
$5.59$ $6698$ $0.997207$ $0.012557$ $1.929368$ $4.355523$
$5.6$ $5082$ $0.996321$ $0.015401$ $1.421225$ $3.490106$
$5.7$ $1501$ $0.987620$ $0.025973$ $1.590960$ $3.434482$
$5.8$ $888.269$ $0.979209$ $0.018833$ $1.305491$ $4.203009$
$5.9$ $635.099$ $0.971088$ $0.021934$ $1.977612$ $4.803605$
66 $496.670$ $0.963240$ $0.027122$ $1.420679$ $3.567866$
$6.1$ $409.353$ $0.955640$ $0.016196$ $1.713398$ $1.710689$
$6.2$ $349.287$ $0.948293$ $0.019949$ $1.623774$ $3.117029$
$6.3$ $305.581$ $0.941204$ $0.066228$ $1.701866$ $3.427829$
$6.4$ $272.026$ $0.934286$ $0.023358$ $1.163604$ $3.125341$
$6.5$ $246.170$ $0.927735$ $0.255900$ $1.454736$ $3.093587$
$6.75$ $199.284$ $0.911787$ $0.057471$ $1.361313$ $4.269028$
77 $168.924$ $0.897081$ $0.059017$ $1.777110$ $3.321819$
88 $109.668$ $0.847438$ $0.038048$ $1.707618$ $2.623961$
99 $84.858$ $0.808831$ $0.031281$ $1.249275$ $2.976333$
$10$ $71.220$ $0.777942$ $0.016334$ $1.624690$ $4.127955$
$11$ $62.611$ $0.752706$ $0.084730$ $1.177847$ $1.792189$
$12$ $56.653$ $0.731621$ $0.020106$ $1.606738$ $2.688990$
$13$ $52.306$ $0.713814$ $0.041793$ $1.671118$ $3.367414$
Table 3: Values for $Q$ in Corollary 1, with corresponding $W$ and parameters $\beta$, $\epsilon_{1}$, $\sigma_{1}$, $\eta$ (all rounded to six decimal places). Each entry is valid for $t\geq 13$.
Theorem 3.
Let $d\leq 1/W_{0}$, $\epsilon_{1}\leq 1/2$, and $\beta<1$ be positive real parameters, subject to (35) and (38). Define $W>W_{0}$ by
$W:=\left(d\beta(1+a_{\epsilon_{1}}(t_{0})^{-1})-d\right)^{-1},$ (30)
where the function $a_{\epsilon_{1}}(t_{0})$ also depends on $\epsilon_{1}$ and is defined in (34).
Then in a region
$\sigma\geq 1-\frac{1}{W\log t},\qquad t\geq t_{0}\geq H$
where $\zeta(\sigma+it)\neq 0$, we have
$\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq Q_{H}\log t,$
where
$Q_{H}=\max\left\{\lambda_{1}\left(\frac{1}{6}+\frac{2\log\log t_{0}}{\log t_{0}}+\frac{\log A_{3}}{\log t_{0}}\right)+\lambda_{2},\,\frac{1}{(d\beta(1+a_{\epsilon_{1}}(t_{0})^{-1})+d)}\right\},$
with
$\lambda_{1}=\frac{16\beta}{(1-\beta)}\left(1-\frac{8d}{2d+\log t_{0}}\right)^{-2}\qquad\text{and}\qquad\lambda_{2}=\frac{1+\beta}{d(1-\beta)},$ (31)
provided that
$t^{1/6}\log t\geq\frac{\zeta(\frac{3}{2}+\frac{2d}{\log t_{0}})}{0.618(1+a_{1}(\frac{3}{2}+\frac{2d}{\log t_{0}},1.31,t_{0}))^{1/6}}$ (32)
is satisfied.
The quantity $A_{3}$ is defined in (4.2), and depends on $d$ and $t_{0}$, on $a_{0}$ and $a_{1}$, which are defined as in Lemma 4, and on $V(\kappa,\sigma_{1},t_{0},\eta)$ and its corresponding parameters, defined as in Lemma 12 and subject to (37).
Proof.
The proof is similar to that of Theorem 2. Suppose $t_{0}\geq H$ and let there be a positive real parameter $d\leq 1/W_{0}$ to be determined later. As before, we construct discs which have origins slightly to the right of the $1$-line, at $s_{0}=\sigma_{0}+iT_{0}$, where $T_{0}\geq t_{0}$,
$\sigma_{0}=1+\frac{d}{\log T_{0}},\qquad r=\frac{1}{2}+\frac{d}{\log T_{0}},$ (33)
extend their radii slightly into the critical strip, and then apply Lemma 2 with $f(s)=\zeta(s)$, $s=\sigma+iT_{0}$, and $A_{1}$ defined by (25). This time however, we do not have the assumption of $RH(T)$ to help us with potential zeros. We will instead utilise the zero-free region mentioned in §2.1 and keep our discs within said region. This requires discs with radii decreasing to 0 at the rate of $1/\log T_{0}$, similar to the zero-free region.
Let $\epsilon_{1}$ be a fixed positive real parameter, and for the sake of brevity define the decreasing function (for $T_{0}>1$)
$a_{\epsilon_{1}}(T_{0}):=1+\frac{\log\left(1+\frac{\epsilon_{1}}{\log T_{0}}\right)}{\log T_{0}}>1.$ (34)
To ensure that the non-vanishing condition on $f(s)=\zeta(s)$ is met, notice that if we impose the condition $\alpha r\leq\epsilon_{1}\leq r$, then the disc $|s-s_{0}|\leq\alpha r$ with
$\alpha r\leq\sigma_{0}-\left(1-\frac{d}{\log(T_{0}+\epsilon_{1})}\right),$
is completely contained in the zero-free region as defined in §2.1. This also implies that the following choice of $\alpha$ and bounds are admissible:
$\alpha:=\frac{2d\left(1+a_{\epsilon_{1}}(t_{0})^{-1}\right)}{\log T_{0}+2d}<\frac{4d}{2d+\log t_{0}}\qquad\text{and}\qquad\frac{2d}{\log t_{0}}\leq\epsilon_{1}\leq\frac{1}{2},$ (35)
which in turn requires that $d<(\log t_{0})/6$ since we need $\alpha<1/2$ for Lemma 2. This is easily achieved since $d\leq 1/W_{0}<(\log t_{0})/6$ for all $t_{0}\geq 3$.
Next, we determine $A_{1}$ and $A_{2}$. Suppose (6) is satisfied with $1+\delta_{r}=\sigma_{0}+r$ and define $A_{1}$ (which depends on $A_{3}$) by (25), but use (33) with the argument of Lemma 12 to replace (4.1) by
$A_{3}:=0.618\left(1+a_{0}\left(\frac{3}{2}+\frac{2d}{\log t_{0}},1.31,t_{0}\right)\right)\left(1+a_{1}\left(\frac{3}{2}+\frac{2d}{\log t_{0}},1.31,t_{0}\right)\right)^{7/6}$
$\bm{\cdot}\left(1+\frac{1}{2t_{0}}+\frac{d}{t_{0}\log t_{0}}\right)^{1/6}\left(1+\frac{\log\left(1+\frac{1}{2t_{0}}+\frac{d}{t_{0}\log t_{0}}\right)}{\log t_{0}}\right)V\left(\frac{d}{\log t_{0}},\sigma_{1},t_{0},\eta\right),$ (36)
where $\sigma_{1}$ and $\eta$ satisfies
$1+\frac{d}{\log t_{0}}\leq\sigma_{1}\leq 2\qquad\text{and}\qquad 0<\eta<2t_{0}.$ (37)
On the other hand, $A_{2}$ is determined by invoking (28), which gives us
$\left|\frac{\zeta^{\prime}(s_{0})}{\zeta(s_{0})}\right|\leq A_{2}:=\frac{1}{d}\log T_{0}.$
Now, for any results from Lemma 2 to be meaningful for us, we require the left most point of $|s-s_{0}|\leq\alpha\beta r$ to lie within the zero-free region and in the critical strip. In other words, $\sigma_{0}-\alpha\beta r\leq 1$, which implies
$\beta\geq\frac{2d}{\alpha(2d+\log T_{0})}\geq\left(1+\frac{1}{a_{\epsilon_{1}}(t_{0})}\right)^{-1}\geq\frac{1}{2}.$ (38)
In particular, since
$\sigma_{0}-\alpha\beta r=1-\left(d\beta(1+a_{\epsilon_{1}}(t_{0})^{-1})-d\right)\frac{1}{\log T_{0}},$
to obtain a result that holds for
$1-\left(d\beta(1+a_{\epsilon_{1}}(t_{0})^{-1})-d\right)\frac{1}{\log T_{0}}\leq\sigma\leq 1+\left(d\beta(1+a_{\epsilon_{1}}(t_{0})^{-1})+d\right)\frac{1}{\log T_{0}},$
we choose $W>W_{0}$ such that
$W:=\left(d\beta(1+a_{\epsilon_{1}}(t_{0})^{-1})-d\right)^{-1},$
while to cover the remaining range of $\sigma\geq 1+\left(d\beta(1+a_{\epsilon_{1}}(t_{0})^{-1})+d\right)/\log T_{0}$, we again use (28), and take the maximum of the two results.
Finally, putting everything together into Lemma 2 and making the substitution $T_{0}\mapsto t$, we prove our result. ∎
4.3 Combining both cases $t\leq H$ and $t\geq H$
Proof of Corollary 1.
For the case of $t\geq H$, with a fixed choice of $W$ as defined in (30), it remains to optimise $Q_{H}$ in Theorem 3 over $d$, $\beta$, $\epsilon_{1}$, $\sigma_{1}$, and $\eta$, subject to the conditions mentioned in said theorem, while also checking that (32) holds, which is done in a similar way as in the proof of Corollary 3. We found that the optimal value for $d$ in all cases is $d=1/W_{0}$, and (32) holds for all $t\geq t_{0}\geq H$. All computed parameters are presented in Table 3, rounded to six decimal places.
On the other hand, a simple calculation shows that for $t\geq t_{0}=13$,
$1-\frac{1}{W_{0}\log t}>0.929.$
Now, for $\sigma\geq 0.9$ and $13\leq t\leq H$, Corollary 3 and Table 2 asserts that
$\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq Q_{RH(H_{0})}\log t=13.123\log t,$
which is smaller than any value of optimised $Q_{H}$ found. Thus this same value of $Q_{H}$ holds for all $t\geq 13$. ∎
We also prove Corollary 2, especially relevant for $\sigma=1$.
Proof of Corollary 2.
First note that Corollary 3 and Table 2 has for $\sigma\geq 1$ and $13\leq t\leq H$,
$\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq Q_{RH(H_{0})}\log t=8.101\log t.$ (39)
Then for $t\geq H$, the proof is identical to that of Theorem 3 and Corollary 1, except we can restrict the size of $\alpha\beta r$ even further, since we can now set $\sigma_{0}-\alpha\beta r=1$, which is equivalent to
$\beta=(1+a_{\epsilon_{1}}(t_{0})^{-1})^{-1}.$ (40)
So in this case, for $\sigma\geq 1$ and $t\geq t_{0}\geq H$, we have that
$\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq\max\left\{\lambda_{1}\left(\frac{1}{6}+\frac{2\log\log t_{0}}{\log t_{0}}+\frac{\log A_{3}}{\log t_{0}}\right)+\lambda_{2},\,\frac{1}{2d}\right\}\,\log t,$ (41)
where $\lambda_{1}$ and $\lambda_{2}$ are defined as in (31), but with the added restriction of (40).
Similar to the proof of Corollary 1, we then optimise (41) over $d$, $\epsilon_{1}$, $\sigma_{1}$, and $\eta$, while ensuring that we fulfil the appropriate conditions in Theorem 3 and (32). The optimal value of $d$ found was again $d=1/W_{0}$. We then compare the result with (39) and take the maximum to obtain a bound that holds for all $t\geq 13$. ∎
5 Estimates for $|1/\zeta(s)|$
Moving from a bound on the logarithmic derivative of $\zeta$ to one on the reciprocal of $\zeta$ is done in the usual way (see [26, Theorem 3.11]). It revolves around estimating the right-hand side of
$\log\left|\frac{1}{\zeta(\sigma+it)}\right|=-\textup{Re}\log\zeta\left(1+\delta_{1}+it\right)+\int_{\sigma}^{1+\delta_{1}}\textup{Re}\left(\frac{\zeta^{\prime}}{\zeta}(x+it)\right)\text{d}x$ (42)
for some $\delta_{1}>0$.
The estimation of the first term on the right-hand side can be improved using the trigonometric polynomial (see [3, Proposition A.2], [20]), while the above integral is estimated by bounding the real part of the logarithmic derivative by a uniform bound on its absolute value $|\zeta^{\prime}(s)/\zeta(s)|$. Here some savings can be made by using the method of [7, Lemma 10]), which splits the interval $\sigma,1+\delta_{1}]$ and uses the sharpest bounds for $|\zeta^{\prime}(s)/\zeta(s)|$ in the appropriate range, rather than one uniform bound for the entire interval.
Furthermore, through an argument utilising the trigonometric polynomial, we are able to get a slight asymptotic improvement on the bounds, at the cost of a worse constant.
Theorem 4.
Let $t_{0}\geq 13$ and $d_{1}>0$ be a constant. Let $\{W_{j}:1\leq j\leq J\}$ be a sequence of increasing real numbers where $(W_{j},Q_{1,j})$ is a pair for which $t\geq t_{0}$ and
$\sigma\geq 1-\frac{1}{W_{j}\log t}\quad\text{implies}\quad\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq Q_{1,j}\log t.$
Then for
$\sigma\geq 1-\frac{1}{W_{1}\log t},$
we have
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq Y_{0}(d_{1},\sigma_{1},t_{0},\eta)\log t,$
where
$Y_{0}$ $(d_{1},\sigma_{1},t_{0},\eta)=\max\Biggl{\{}\frac{1}{d_{1}}+\frac{1}{\log t_{0}},$ (43)
$V(d_{1}/\log t_{0},\sigma_{1},t_{0},\eta)\cdot\exp\Bigg{(}\sum_{j=1}^{J-1}Q_{1,j}\left(\frac{1}{W_{j}}-\frac{1}{W_{j+1}}\right)+\frac{Q_{1,J}}{W_{J}}+24.303d_{1}\Bigg{)}\Biggr{\}},$
with any $1+d_{1}/\log t_{0}\leq\sigma_{1}\leq 2$, $\gamma$ denoting Euler’s constant, and $V$ defined as in Lemma 12.
Proof.
Let $\delta_{1}=\delta_{1}(t):=d_{1}/\log t$. We will split the proof into two cases: $\sigma\leq 1+\delta_{1}$ and $\sigma>1+\delta_{1}$. First see that in the range $1-1/W_{1}\log t\leq\sigma\leq 1+\delta_{1}$ we have
$\log\left|\frac{1}{\zeta(\sigma+it)}\right|$ $=-\textup{Re}\log\zeta\left(1+\delta_{1}+it\right)+\int_{\sigma}^{1+\delta_{1}}\textup{Re}\left(\frac{\zeta^{\prime}}{\zeta}(x+it)\right)\text{d}x.$ (44)
Writing $\Delta=1/\log t$, we can split the integral and rewrite it as
$\left(\int_{1-\tfrac{\Delta}{W_{1}}}^{1-\tfrac{\Delta}{W_{2}}}+\cdots+\int_{1-\tfrac{\Delta}{W_{j}}}^{1-\tfrac{\Delta}{W_{j+1}}}+\cdots+\int_{1-\tfrac{\Delta}{W_{J}}}^{1}+\int_{1}^{1+\delta_{1}}\right)\textup{Re}\left(\frac{\zeta^{\prime}}{\zeta}(x+it)\right)\text{d}x.$ (45)
By Corollary 2 and our assumptions we have
$\int_{\sigma}^{1+\delta_{1}}\textup{Re}\left(\frac{\zeta^{\prime}}{\zeta}(x+it)\right)\text{d}x$ $\leq\sum_{j=1}^{J-1}Q_{1,j}\left(\frac{1}{W_{j}}-\frac{1}{W_{j+1}}\right)+\frac{Q_{1,J}}{W_{J}}+24.303d_{1}.$
We therefore have, from (44),
$\log\left|\frac{1}{\zeta(\sigma+it)}\right|<$ $-\log\left|\zeta\left(1+\frac{d_{1}}{\log t}+it\right)\right|$ (46)
$+\sum_{j=1}^{J-1}Q_{1,j}\left(\frac{1}{W_{j}}-\frac{1}{W_{j+1}}\right)+\frac{Q_{1,J}}{W_{J}}+24.303d_{1}.$
To estimate the first term on the right-hand side of (46) we apply Lemma 12, keeping in mind our restrictions on $\sigma_{1}$. This gives us
$\left|\frac{1}{\zeta(1+\delta_{1}+it)}\right|\leq V(d_{1}/\log t_{0},\sigma_{1},t_{0},\eta)\log t.$ (47)
We combine this with (46) and exponentiate to obtain $Y_{0}$ for $\sigma\leq 1+\delta_{1}$.
For the case $\sigma>1+\delta_{1}$ we have from the trivial bound $|1/\zeta(s)|\leq\zeta(\sigma)$ with (9) that
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq\left(\frac{1}{d_{1}}+\frac{1}{\log t_{0}}\right)\log t,$ (48)
since $\sigma/(\sigma-1)$ is decreasing in $\sigma$.
Taking the maximum from both cases completes the proof. ∎
Corollary 4 follows from Theorem 4 by computing $Y_{0}$ for specific choices of $t_{0}$ and $W_{j}$, and optimizing over $d_{1}$, $\sigma_{1}$, and $\eta$. These computed values of $Y_{0}$ are labelled $Y$, and are presented in Table 4, with its corresponding parameters (rounded to six decimal places). The process is the same as that of [7, Corollary 3], but for completeness sake, we repeat it here.
Proof of Corollary 4.
After choosing $t_{0}=13$, we aim to minimise $Y_{0}$ by optimising over $d_{1}>0$, $\eta$, and $\sigma_{1}$, while making sure that the conditions on these parameters are fulfilled.
This optimisation process first requires fixing the desired $W_{1}$, which then tells us the number of $W_{j}$ we are summing over, since from Table 3 we always have $(W_{J},Q_{1,J})=(13,52.306)$. Computing more values of $(W_{J},Q_{1,J})$ will give a better result, but the improvements eventually become negligible, as $Y_{0}$ is largely determined by the initial few $W_{j}$. For our computations, we only used values for $W_{j}=W$ from Table 3 (Corollary 1), thus we require $t_{0}=13$.
Note that one can adjust Theorem 4 to compute $Y$ for the case $\sigma\geq 1$. In that case, we will modify (43) by setting
$\sum_{j=1}^{J-1}Q_{1,j}\left(\frac{1}{W_{j}}-\frac{1}{W_{j+1}}\right)+\frac{Q_{1,J}}{W_{J}}=0$
before optimising. We obtain $Y=30.812$ for $\sigma\geq 1$ and $t\geq 13$, with $(d_{1},\sigma_{1},\eta)=(0.032871,1.662479,3.216997)$. This proves the second assertion of the corollary.
To prove the third assertion, we first compute $Y=29.388$ for $\sigma\geq 1$ and $t\geq 2\exp(e^{2})$, with $(d_{1},\sigma_{1},\eta)=(0.034172,1.286253,3.718165)$. Then see that from Lemma 7,
$\left|\frac{1}{\zeta(1+it)}\right|\leq\frac{8}{\log(500)}\log t\leq 1.288\log t\quad\text{for}\quad 500\leq t\leq 2\exp(e^{2}),$
while a computation from the proof of [3, Proposition A.2], implemented in interval arithmetic, states that
$\left|\frac{1}{\zeta(1+it)}\right|\leq 2.079\log t\quad\text{for}\quad 2\leq t\leq 500,$ (49)
so we are done. ∎
$W$ $Y$ $d_{1}$ $\sigma_{1}$ $\eta$
$5.56$ $7.59\cdot 10^{31}$ $0.030648$ $1.151072$ $3.150608$
66 $2.46\cdot 10^{9}$ $0.030648$ $1.150464$ $3.150009$
77 $1.83\cdot 10^{6}$ $0.030647$ $1.151694$ $3.149136$
88 $89357$ $0.030647$ $1.148849$ $3.149156$
99 $19482$ $0.030647$ $1.150203$ $3.150832$
$10$ $7589$ $0.030648$ $1.149868$ $3.150556$
$11$ $3972$ $0.030650$ $1.150918$ $3.151137$
$12$ $2472$ $0.030646$ $1.148656$ $3.150041$
$13$ $1719$ $0.030647$ $1.149567$ $3.150198$
Table 4: Values for $Y$ in Corollary 4, with corresponding $W$, and parameters (rounded to six decimal places): $d_{1}$, $\sigma_{1}$, and $\eta$. Each entry is valid for $t\geq 13$.
5.1 An asymptotic improvement
In this section we prove an asymptotic improvement over the bounds given earlier in §5, at the expense of a worse constant. For the sake of clarity, we first outline the general idea here, before proving the next theorem.
Suppose we have the bound
$\max_{1-\delta\leq\sigma\leq 1+\delta_{1}}\left|{\textup{Re}}\left(\frac{\zeta^{\prime}}{\zeta}(\sigma+it)\right)\right|\leq C\log t,$
for $t\geq t_{0}$ and $1-\delta\leq\sigma\leq 1+\delta_{1}$, where $\delta=d/\log t$ and $\delta_{1}=d_{1}/\log t$, as before. Then from (42), we have
$\left|\frac{1}{\zeta(\sigma+it)}\right|$ $\leq\left|\frac{1}{\zeta(1+\delta_{1}+it)}\right|\exp\left((\delta_{1}+\delta)\max_{1-\delta\leq\sigma\leq 1+\delta_{1}}\left|{\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(\sigma+it)\right|\right)$
$\leq\left|\frac{1}{\zeta(1+\delta_{1}+it)}\right|\exp\left((d_{1}+d)C\right).$ (50)
Notice that since the exponential factor is $O(1)$, the overall order is determined by the first factor.
With the same argument involving the trigonometric inequality (20),
$\left|\frac{1}{\zeta(1+\delta_{1}+it)}\right|\leq|\zeta(1+\delta_{1})|^{3/4}|\zeta(1+\delta_{1}+2it)|^{1/4},$ (51)
and at the expense of a worse constant, we can estimate the second factor in (51) differently from Theorem 4. We do this by using Lemma 6, which states that $|\zeta(s)|\leq C_{3}(\log t)^{2/3}$ for $\sigma\geq 1$. Along with the second inequality in Lemma 9 and our choice of $\delta_{1}$, this leads to
$\left|\frac{1}{\zeta(1+\delta_{1}+it)}\right|\leq(C_{4}\log t)^{3/4}(C_{3}(\log t)^{2/3})^{1/4},$
which applied to (5.1) ultimately furnishes the bound
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq C_{5}(\log t)^{11/12}.$
Theorem 5.
Let the hypotheses of Theorem 4 be satisfied. Then for $t\geq t_{0}\geq 13$ and
$\sigma\geq 1-\frac{1}{W_{1}\log t},$
we have
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq Y^{\prime}_{0}(d_{1},t_{0})(\log t)^{11/12},$
where
$Y^{\prime}_{0}$ $(d_{1},t_{0})=C_{3}^{1/4}\cdot\max\Bigg{\{}\left(\frac{1}{d_{1}}+\frac{1}{\log t_{0}}\right)^{3/4},$
$\exp\Bigg{(}\sum_{j=1}^{J-1}Q_{1,j}\left(\frac{1}{W_{j}}-\frac{1}{W_{j+1}}\right)+\frac{Q_{1,J}}{W_{J}}+24.303d_{1}\Bigg{)}\left(\frac{e^{\gamma d_{1}/\log t_{0}}}{d_{1}}\right)^{3/4}\Bigg{\}},$
with $d_{1}$, $J$, $Q_{1,j}$, $W_{j}$ defined as in Theorem 4, and $C_{3}$ defined as in Lemma 6.
Proof.
The proof is identical to that of Theorem 4, except that in the case $\sigma\leq 1+\delta_{1}$, we replace (47) with
$\left|\frac{1}{\zeta(1+\delta_{1}+it)}\right|$ $\leq|\zeta(1+\delta_{1})|^{3/4}|\zeta(1+\delta_{1}+2it)|^{1/4}$
$\leq\left(\frac{e^{\gamma(d_{1}/\log t_{0})}}{d_{1}}\log t\right)^{3/4}\left(C_{3}(\log t)^{2/3}\right)^{1/4}$
by using (51) with Lemmas 6 and 9.
Meanwhile, for $\sigma>1+\delta_{1}$, the same argument with (20) and (9) leads to
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq\left(\left(\frac{1}{d_{1}}+\frac{1}{\log t_{0}}\right)\log t\right)^{3/4}\left(C_{3}(\log t)^{2/3}\right)^{1/4}$
in place of (48). ∎
Proof of Corollary 5.
The proof, and process of optimisation here, even for the case $\sigma\geq 1$, are exactly the same as that of Corollary 4. These computed values of $Y^{\prime}_{0}$ for $t\geq 13$, are labelled $Y^{\prime}$, and are presented in Table 5. For all these cases, we found an optimal parameter of $d_{1}=0.030648$ (rounded to six decimal places).
For the case $\sigma\geq 1$, we computed $Y^{\prime}=87.725$ for $t\geq 13$, also with optimal $d_{1}=0.030648$ (rounded).
Finally, for the case $\sigma=1$, we computed $Y^{\prime}=81.900$ for $\sigma\geq 1$ and $t\geq 2\exp(e^{2})$, with $d_{1}=0.030793$ (rounded). Then by a similar calculation as in the proof of Corollary 4, we have by Lemma 7,
$\left|\frac{1}{\zeta(1+it)}\right|\leq 1.5(\log t)^{11/12}\quad\text{for}\quad 500\leq t\leq 2\exp(e^{2}).$
On the other hand, (49) implies
$\left|\frac{1}{\zeta(1+it)}\right|\leq 2.421(\log t)^{11/12}\quad\text{for}\quad 2\leq t\leq 500,$
where this estimate is obtained by noting that $t$ attains its maximum at $500$. ∎
$W$ $Y^{\prime}$
$5.56$ $2.17\cdot 10^{32}$
66 $6.99\cdot 10^{9}$
77 $5.21\cdot 10^{6}$
88 $2.55\cdot 10^{5}$
99 $55579$
$10$ $21649$
$11$ $11331$
$12$ $7051$
$13$ $4904$
Table 5: Values for $Y^{\prime}$ in Corollary 5, with corresponding $W$. Each entry is valid for $t\geq 13$.
6 Concluding Remarks
Here we state some possible ways to improve our estimates. The most impactful would be having a larger zero-free region or a higher Riemann height. Using an asymptotically stronger zero-free region is also helpful. See for instance [7], where a Littlewood-style zero-free region is used to obtain bounds of order $\log t/\log\log t$. Minor improvements can be made by a more careful treatment of the sum over zeros and the integral arising from the application of Lemma 2. For Theorem 4, a better smoothing argument could be used in place of (45).
Other possibilities for improvement are: a refinement on the Borel–Carathéodory by specialising it more to this problem, using stronger estimates for $\zeta(s)$ under $RH$ or $RH(T)$, or obtaining a good bound on ${\textup{Re}}(\zeta^{\prime}(s)/\zeta(s))$ for use in (42).
It is evident from the results that the closer $s$ is to the boundary of the zero-free region, and hence a potential zero, the larger the constants. However, the large constants can also be attributed in part to the factor (R+r)/(Rr)(R+r)/(R-r) in Theorem 1, which increases rapidly as $r$ approaches $R$. This contributes significantly to the overall estimate, even if $|f(s_{0})|$ is small, and seems to suggest that Theorem 1 in its current form might not be optimal for bounding $|\zeta^{\prime}(s)/\zeta(s)|$.
An advantage of obtaining a bound on ${\textup{Re}}(\zeta^{\prime}(s)/\zeta(s))$ is to reduce the usage of Theorem 1, when bounding $|1/\zeta(s)|$. This is because we would be able to apply such a bound directly to (42), whereas we currently use ${\textup{Re}}(\zeta^{\prime}(s)/\zeta(s))\leq|\zeta^{\prime}(s)/\zeta(s)|$. This alternative could be approached via Lemma 2, provided one were able to deal with the resulting sum over zeros well.
To get unconditional results we have (due to knowledge of the Riemann height) assumed a partial Riemann hypothesis to prove some of our results. The downside to this is that we have not used better available bounds that require the full strength of $RH$. For example, under $RH$ we could have used that $\zeta(1/2+it)=O(\exp(C\log t/\log\log t))$ for some $C>0$ to prove something stronger (see [26, §14], [25], [24], and [6] for more details).
Even in the unconditional results used, there is room for improvement. As an illustration, Corollary 5, though asymptotically superior, beats Corollary 4 only when $t$ is large. In the case of $\sigma=1$, this happens for $t\geq\exp(2.2\cdot 10^{5})$. This is in part due to the relatively large constants in Corollary 5, and a direct way of lowering these constants is an improvement in the bound $|\zeta(1+it)|$, which correspondingly lowers $C_{3}$ in Lemma 6.
Finally, we conclude with a discussion on the trigonometric inequality (20) used in many of our above proofs, in particular when bounding $|1/\zeta(s)|$ for $\sigma>1$. Under certain conditions, this method is superior to using the trivial bound. In general, when studying zero-free regions for $\zeta(s)$, increasing the degree of the non-negative trigonometric polynomial improves the results one gets. Surprisingly, the same cannot be said in our setting, and the classical polynomial we used is actually optimal. Take for instance Lemma 12, which is used in Theorem 4. Here the overall order of the bound is $O(\log t)$, and in this scenario the article [20] has an in-depth discussion on why the classical polynomial is best.
On the other hand, in Theorem 5, the question is whether higher degree polynomials lead to a reduction in the exponent $11/12$, even if overall constants get worse. This is an interesting question to consider, since in this case the answer is less obvious, although preliminary calculations suggest that the classical polynomial is still the optimal one.
Acknowledgements
We thank Tim Trudgian for his guidance throughout the writing of this paper. We also thank the referee for the care taken in going through our article. Finally, we are very grateful to Andrew Fiori, Bryce Kerr, Olivier Ramaré, Aleksander Simonič, for the helpful discussions, and especially to Harald Helfgott for pointing out an error in an earlier version of this paper.